### v2

# Requirements-Based Test Generation: A Comprehensive Survey

Abstract.
As an important way of assuring software quality, software testing generates and executes test cases to identify software failures. Many strategies have been proposed to guide test-case generation, such as source-code-based approaches and methods based on bug reports. Requirements-based test generation (RBTG) constructs test cases based on specified requirements, aligning with user needs and expectations, without requiring access to the source code. Since its introduction in 1994, there have been many contributions to the development of RBTG, including various approaches, implementations, tools, assessment and evaluation methods, and applications. This paper provides a comprehensive survey on RBTG, categorizing requirement types, classifying approaches, investigating types of test cases, summarizing available tools, and analyzing experimental evaluations. This paper also summarizes the domains and industrial applications of RBTG, and discusses some open research challenges and potential future work.
Software testing, test generation, software requirements, survey
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Requirements analysisccs: General and reference Surveys and overviews
1. Introduction
Software development is a complex process driven by rapidly evolving technologies and continuously changing requirements. It is often influenced by shifts in business processes. A key part of the software-development process is software testing, which helps to ensure the quality of the software. Testing involves verifying that the software satisfies the specified requirements, and can account for over half of the total development costs (Mustafa et al., 2021). There are three main stages in software testing: test case generation, execution, and evaluation. Among these, test-case generation may represent the most intellectually demanding process. Manual test-case generation can be time-consuming and prone to human error.
Automated test-case generation has been widely regarded as an effective approach to reducing effort and cost (Jin and Lano, 2021). Test cases can be derived from requirements specifications, design artifacts (Zakeriyan et al., 2021), source code (Hamberger et al., 2023), or historical test artifacts (such as bug reports) (Plein et al., 2024). While code-based test-case generation can effectively identify errors at the implementation level, it may not address issues arising from misunderstandings or mistakes in the analysis or design phases (Yazdani Seqerloo et al., 2019). Faulty code can also mislead the test-generation process. Automating test-case generation from bug reports involves reproducing reported issues following the steps provided by users (Plein et al., 2024; Fazzini et al., 2018). This can be valuable for managing and analyzing a large volume of bug reports, thus reducing the manual effort required. However, there may be challenges, due to the informal nature of user-reported bugs, and that these reports may be mainly from the software maintenance phase. Requirements-based test generation (RBTG), which relies on requirements specifications or models (before any code is written), may provide a more robust solution. By generating test cases early in the development lifecycle, RBTG can detect inconsistencies and ambiguities in requirements, reducing the risk of propagating errors into later stages of development. This early intervention can significantly lower development costs by addressing defects earlier.
Due to its simplicity, and no specialized training being needed to understand it, natural language (NL) has often been used for software requirements’ documentation (Dwarakanath and Sengupta, 2012). However, NL can be imprecise and ambiguous, potentially making it unsuitable for automated test-case generation. Over the years, significant research has been dedicated to addressing the challenge of RBTG (Rodrigues et al., 2024). To address the limitations of NL, semi-formal and formal representations have also been proposed (Fischbach et al., 2019). RBTG can be traced back to Peng Lu in 1994 (Lu, 1994). Mustafa et al. (2021) presented a review of RBTG from 2000 to 2018, reporting that most approaches used Unified Modeling Language (UML) for RBTG, with a primary focus on functional testing. Farooq and Tahreem (2022), Jin and Lano (2021), and Ahmad et al. (2019) also examined UML-based methodologies for test case generation. Recent studies have also explored the application of natural language processing (NLP) techniques to the testing of NL-based requirements (Boukhlif et al., 2024) (Garousi et al., 2020). Wang et al. (2024) provided a comprehensive analysis of large language models (LLMs) in software testing, while Fontes and Gay (2023) investigated the integration of machine learning (ML) techniques for automated test generation. These studies focused on the application of NLP, LLMs, and ML in test case generation. Nevertheless, to the best of our knowledge, there is a lack of an up-to-date, comprehensive survey encompassing the entire RBTG process and potential future research directions. This paper aims to address this gap.
This paper presents a comprehensive survey of RBTG, covering 267 papers published up to 2024. It examines the entire RBTG process, including: (1) a summary and analysis of the selected 267 papers; (2) a classification of the RBTG input requirements; (3) a description, classification, and synthesis of RBTG approaches; (4) a description of the types of test cases produced; (5) a summary of RBTG tools; (6) an analysis of RBTG evaluation methods; (7) an overview of the RBTG application domains; and (8) a discussion of open research challenges and potential future directions for RBTG. To the best of our knowledge, this is the first large-scale and comprehensive survey on RBTG.
The rest of this paper is structured as follows. Section 2 discusses the background and previous surveys. Section 3 explains the methodology adopted for our literature review. Section 4 explores the distribution of RBTG studies. Section 5 examines the types of RBTG input requirements. Section 6 reviews the approaches used for RBTG. Section 7 categorizes the types of test cases generated. Section 8 presents the tools developed for RBTG. Section 9 provides a detailed analysis of RBTG evaluation methods. Section 10 examines the application domains covered in RBTG studies. Section 11 identifies key RBTG challenges, and proposes future research directions. Finally, Section 12 concludes the paper.
2. Background and Related Work
This section introduces the definition and classification of requirements, and explains their importance for defining system objectives. We also outline the types of tests, explaining how they help to ensure that the system performs as intended. Previous survey studies on test-case generation are also reviewed, summarizing their key outcomes and methodologies.
2.1. Requirements
Software requirements serve as the blueprint for the development of a software application, presenting the core goals and objectives that guide the development team (Umar and Lano, 2024). The requirements specify the intended features and functionality, covering aspects such as product perspective, user characteristics, operating environment, design constraints, user documentation, assumptions, and dependencies (Bruel et al., 2021). The quality of the requirements can significantly impact the overall quality of both the software testing and development (Mustafa et al., 2021).
In typical industrial environments, users may communicate the needs of a software application by stating the requirements in NL, using things like requirements specifications and user stories. Pure NL is easy to understand, making it accessible to both developers and testers. However, requirements described in NL may contain ambiguities or inconsistencies, and may be incomplete. This will impact on the quality of software testing. To minimize such problems, requirements can be converted into more structured and formal representations, enhancing clarity and consistency.
Fischbach et al. (2019) noted that requirements could be specified into three different classes: informal, semi-formal, and formal techniques. Informal descriptions do not follow a particular syntax or semantic structure; semi-formal requirements are expressed in a predefined and fixed syntax that is not formally defined; and formal notations are based on mathematical notation, and have precise syntax and semantics. However, it is difficult to fully represent the diverse forms of requirements within RBTG using just this classification. Furthermore, these broad categories lack the precision needed to align with the trends in modern requirements engineering.
Bruel et al. (2021) categorized requirements into five groups based on their primary specification style: NL, semi-formal, automata/graphs, mathematical, and seamless (programming language-based). Semi-formal specifications structure the expression of requirements by establishing a defined language that is more precise than NL, yet does not rely on mathematical formalisms or programming-language constructs. Examples of this kind of requirement include: Reqtify (Systems, 2016), KAOS (Van Lamsweerde, 2001), and systems modeling language (SysML) (Chen et al., 2023). Automata/graphs-based requirements are modeled through state machines (Shaheen et al., 2024), transition systems (Cartaxo et al., 2007), or graph-theoretical representations, such as UML activity diagrams (ADs) (Kamonsantiroj et al., 2019) and state diagrams (Chevalley and Thevenod-Fosse, 2001). Mathematical notation relies on formalisms based on set theory (Kadakolmath and Ramu, 2022; Saiki et al., 2021) or universal algebra (Dick and Faivre, 1993). Seamless requirements are closely integrated with implementation artifacts through programming language constructs, enabling direct traceability to the code (Naumchev, 2019). In the context of an RBTG review study, the generation of test cases is independent of source code: The “seamless” category, therefore, is problematic, as studies involving source code are typically excluded during the literature search. Nevertheless, within the field of test-case generation methods, model-based approaches are extensively employed, with a wide variety of model types. The “automata/graphs” category outlined in this paper does not comprehensively cover all model types, nor does it account for scenarios involving the combined use of multiple requirements’ types. To address these limitations, this paper proposes a revised classification framework that builds on the existing categories while aligning more closely with the characteristics of RBTG methods.
Table 1. Requirements Classification for RBTG
Category Core Characteristics Typical Formats
Natural language Unstructured or simplified textual descriptions in human language. Unstructured natural language (Aoyama et al., 2021), use case textual description (Wang et al., 2020), user stories (Fischbach et al., 2020)
Semi-formal Combines natural language with structured elements such as templates, tables, or simplified graphical representations Ontology-based requirements specification (Banerjee et al., 2021; Ul Haq and Qamar, 2019), SysML (Zhu et al., 2021; Chen et al., 2023)
Model-based Abstract representations of software behavior or functionality, using graphical, automata, or trees. UML diagrams (Linzhang et al., 2004; Sahoo et al., 2021), Automata (Hopcroft et al., 2001; Linz and Rodger, 2022)
Formal mathematical Mathematically rigorous and precise, using formal notations for specification. Structured Object-Oriented Formal Language (SOFL) (Saiki et al., 2021; Cajica et al., 2021), Z formal specification (Kadakolmath and Ramu, 2022; Helke et al., 1997)
Hybrid Combine requirements from multiple categories for flexibility and comprehensiveness. Textual specifications combined with a use case diagram, textual use cases, HTML screen mockups, and glossaries (Clerissi et al., 2017)
Table 1 groups the requirements representations in the surveyed papers into five categories:
  • Natural language requirements: Expressed in English or other natural languages, these requirements are intuitive and easy to understand. To improve clarity and accessibility for both technical and non-technical audiences, their expressiveness is sometimes restricted to a controlled subset of the language.
  • Semi-formal requirements: Combining NL with some formal structure, such as tables in text documents, or diagrams from modeling tools. These approaches provide enhanced clarity and accessibility for users with minimal technical training, but cannot be directly processed by formal verification methods.
  • Model-based requirements: Representing software behavior or functionality through abstract representations, such as graphical models, automata, or trees. These representations improve understanding and communication for both technical and non-technical users by using visual aids without necessitating deep knowledge of the underlying mathematics. These models (e.g., UML-based models) have a long tradition in software engineering (Uzun and Tekinerdogan, 2018).
  • Formal mathematical requirements: Beyond those of the previous categories, these requirements are specified using rigorous mathematical notations (such as those based on set theory or algebra), allowing precise definitions suitable for formal verification. However, these are typically only accessible to technical audiences with a strong foundation in mathematics, making them less appropriate for non-technical stakeholders.
  • Hybrid requirements: Combining requirements from two or more categories, hybrid requirements combine the strengths of each to provide a more comprehensive and detailed representation. This approach ensures both breadth and accuracy in capturing requirements information, making it adaptable to diverse project needs.
2.2. Types of Tests
There are different types of tests, such as abstract test cases (ATCs), concrete test cases (CTCs), and test scenarios (TSs). A test case specifies a particular condition to be evaluated, and includes a set of test inputs, execution conditions, and expected results developed to verify a specific aspect of system behavior. A test case may consist of a title, precondition, test steps, and expected result. The definitions used in this paper conform to standard terminology (Board, 2014).
  • Abstract test cases are high-level test cases without concrete (implementation level) values for input data and expected results. Logical operators are used; instances of the actual values are not yet defined and/or available (e.g., given an abstract value “$n$ is a positive integer”). It is typically in a human-readable description.
  • Concrete test cases are low-level test cases with concrete (implementation level) values for input data and expected results. Logical operators from high-level test cases are replaced by actual values that correspond to the objectives of the logical operators (e.g., a concrete value for $n$ is $100$).
  • Test scenarios are documents specifying a sequence of actions for the execution of a test. These consist of test-design specifications, test-case specifications, and/or test-procedure specifications (semi-formal specifications that lie between descriptive system requirements and executable test cases). Test scenarios have been treated as ATCs in several studies (Sabbaghi, 2021).
In software testing, CTCs can either be executed manually (for example, by test engineers) or automatically, using testing frameworks such as the Selenium framework (Waitchasarn et al., 2023) or JUnit (Uzun and Tekinerdogan, 2018). However, some of the reviewed studies did not actually generate test cases: This paper considered those test artifacts lacking concrete input values or expected results as ATCs.
2.3. Previous Survey Studies
Previous work has explored the challenge of generating test cases directly from software requirements, addressing the problem using a wide range of methods and technologies. This paper provides a comprehensive review of these efforts, synthesizing the existing body of knowledge and categorizing the diverse strategies employed in RBTG. Table 2 provides a comparison of prior survey studies against this paper.
Table 2. Comparison of Previous Survey Studies
Year Studies # Studies Time Span Description
2011 An overview on test generation from functional requirements (Escalona et al., 2011) 24 1988-2009 This survey examined approaches for describing functional requirements informally and formalized the comparative study through characterization schema.
2017 A comprehensive investigation of natural language processing techniques and tools to generate automated test cases (Ahsan et al., 2017) 16 2005–2016 This survey explored NLP techniques and tools for automating test case generation from initial requirements documents.
2018 Automated Regression Test Case Generation for Web Application: A Survey (Gupta et al., 2018) 79 1988-2017 This survey focused on the generation of regression test cases and the approaches to web applications.
2019 Model-based testing using UML activity diagrams: A systematic mapping study (Ahmad et al., 2019) 41 2004-2016 This survey provided a comprehensive overview of existing approaches to model-based testing using UML ADs.
2021 Automated Test Case Generation from Requirements: A Systematic Literature Review (Mustafa et al., 2021) 29 2000-2018 This survey focused on functional testing, examining approaches and input requirement types for generating test cases.
2021 Generation of Test Cases from UML Diagrams - A Systematic Literature Review (Jin and Lano, 2021) 62 1999-2019 This survey investigated the generation of test cases from UML diagrams, focusing on key aspects such as model types, intermediate formats, case studies, coverage criteria, execution methods, and test case optimization.
2022 Requirement-Based Automated Test Case Generation: Systematic Literature Review (Farooq and Tahreem, 2022) 30 2010-2021 This survey evaluated techniques for automated test case generation using requirements as input. It primarily focused on UML diagrams and identified the challenges and gaps.
2023 Current Trends in Automated Test Case Generation (Potuzak and Lipka, 2023) 67 2000-2022 This survey addressed automated test data generation and tests based on it, mapping and categorizing existing methods while summarizing their common features.
2023 Software Test Case Generation Tools and Techniques: A Review (Verma et al., 2023) 125 2011-2021 This survey conducted a systematic study of testing tools and techniques without being limited to those based on requirements.
2024 Natural Language Processing-Based Software Testing: A Systematic Literature Review (Boukhlif et al., 2024) 24 2013-2023 This survey provided an in-depth analysis of NLP-based software testing, exploring NLP techniques, software testing types, key challenges, and insights into the application of NLP in software testing.
2024 Software Testing With Large Language Models: Survey, Landscape, and Vision (Wang et al., 2024) 104 2020-2023 This paper presented a detailed discussion of software testing tasks where LLMs are frequently used, with test case preparation and program repair being the most prominent.
2025 This survey 267 1994-2024 This survey investigated the entire process of RBTG, including input representations, generation approaches, supporting tools, evaluation methods, and application domains. It also highlighted open research challenges and proposed potential future directions.
Boukhlif et al. (2024) presented an overview of diverse NLP techniques used to address various linguistic challenges. They provided an analysis of the distribution of software testing contexts across the reviewed research papers, noting a wide range of testing methodologies and objectives. They noted the critical importance of both machine- and deep-learning techniques in NLP for enhancing the capabilities and performance of language-centric applications. However, their study did not specify which techniques were best suited for particular testing methodologies. Nor did they explain how these integrated NLP and machine/deep learning approaches were applied to advance software testing practices.
Mustafa et al. (2021) focused on functional testing, presenting input requirements and the approaches used for requirements transformation. However, they did not establish a clear connection between the requirements types and the corresponding methods, leaving a gap in how these elements are aligned. Similarly, Farooq and Tahreem (2022) explored the generation of functional test cases, with particular attention on the use of UML diagrams: They identified several gaps and challenges in RBTG, offering insights into areas requiring further investigation.
Ahmad et al. (2019) provided a comprehensive overview of model-based testing using UML ADs. They note that only a limited number of their surveyed approaches have been validated against realistic, industrial case studies, with most focusing on very restricted application domains. They also report that UML ADs are not commonly employed for non-functional testing, and that there is a lack of holistic or comprehensive approaches for model-based testing using UML ADs.
Previous reviews primarily focused on input requirements and detailing the approaches and tools utilized in RBTG. However, these studies lack a comprehensive analysis of the interconnections between requirements, methods, tools, and processes within the context of the complete RBTG workflow. To address this gap, this paper reports on a systematic literature review (SLR): It seeks to overcome limitations of previous work, and to provide deeper insights into the different types of requirements, and their corresponding methodologies, tools, and evaluation frameworks. The paper examines the complete RBTG workflow.
3. Research Methodology
We conducted a comprehensive investigation of RBTG using the SLR methodology proposed by Kitchenham et al. (2009). As part of the SLR, we adopted the snowballing technique, which starts with an initial set of papers and expands by examining their references to identify additional relevant studies (Wohlin, 2014). The research methodology consisted of several key steps: defining research questions; identifying relevant studies; applying the snowballing technique to expand the search; implementing exclusion criteria; and systematically extracting the required data. This structured approach ensures a rigorous and comprehensive analysis of the literature.
3.1. Research Questions
Figure 1. Requirements-based test-generation process.
[…]
Requirements-based test-generation process.
The research objectives for this survey were based on the complete RBTG process (Figure 1), and involved analyzing and categorizing information at each stage. The study also aimed to identify research challenges and areas requiring future work. To achieve these objectives, we defined the following research questions (RQs):
  • RQ1: What are the publication patterns of RBTG studies?
  • RQ2: What types of requirements have been used in RBTG?
  • RQ3: What are the approaches that support RBTG?
  • RQ4: What types of test-case representations have been explored in RBTG?
  • RQ5: What tools have been developed to support RBTG?
  • RQ6: How have RBTG approaches been evaluated?
  • RQ7: In what domains and industrial applications has RBTG been applied?
  • RQ8: What are the remaining challenges and future work for RBTG?
3.2. Literature Search and Selection
The search strategy was optimized to identify relevant information, ensuring a thorough and effective investigation of the research questions. This process involved four steps: selecting appropriate digital libraries; identifying additional search sources; determining relevant search strings; and performing the snowballing technique. Following previous review studies (Ahmad et al., 2019; Jin and Lano, 2021; Mustafa et al., 2021; Verma et al., 2023), we selected five online source repositories to perform the SLR:
  • ACM Digital Library
  • Elsevier Science Direct
  • IEEE Xplore Digital Library
  • Springer Online Library
  • Wiley Online Library
These selected repositories provide a wide range of scholarly resources highly relevant to the research topic.
After identifying the relevant literature databases, a search string was applied to each. The search strings were developed iteratively, being refined based on preliminary results. For each repository, customized search strings were constructed to optimize the advanced search process. The overall strategy for constructing the strings was: (1) The Title should include test case OR test suite OR test scenario, (2) The Title should also include generate OR create OR requirement, (3) The Abstract should include requirement OR specification OR user stories. These three criteria were combined using the AND operator to construct the search strings.
The search was conducted in January 2025, capturing all studies published before that date. Table 3 shows the initial search results: These were then de-duplicated and screened based on titles and abstracts, as reported in the third column. To further enhance our literature search, we employed Snowballing techniques (Wohlin, 2014) were used to enhance the literature search: The references of selected papers were reviewed, and additional relevant studies were identified (and downloaded from Google Scholar). This iterative process was continued until no new relevant papers were identified. The final results are listed in the “Post-Snowballing” column. The final set of relevant papers was determined by applying the following exclusion criteria:
  1. (1)
    Studies without a full text.
  2. (2)
    Studies not written in English.
  3. (3)
    Books, book chapters, or technical reports (most of which may have been published as articles).
  4. (4)
    Masters or Ph.D. theses.
  5. (5)
    Survey studies.
  6. (6)
    Studies less than five pages in length.
Table 3. Overview of Search Result and Study Selection
Source
Initial Search
Results
After De-duplication &
Title/Abstract Review
Post-Snowballing
Final After
Exclusions
ACM 76 30 49 31
ELSEVIER 46 12 19 14
IEEE 314 127 182 113
Springer 34 6 57 51
Wiley 23 4 6 3
Other 55 55
Total 493 179 368 267
Table 3 also shows the distribution of selected studies across the various digital libraries: 45% from IEEE Xplore; 18% from SpringerLink; 12% from ACM Digital Library; 5% from Elsevier; 1% from Wiley; and 19% from other databases (such as ProQuest, Academia.edu, and SCIRP.org), identified through snowballing techniques. This diverse distribution highlights the collaborative and extensive nature of academic research in the field.
Although it may not have been possible to find every relevant paper through our search, we are confident that we have included most of the significant publications. This survey provides a comprehensive overview of the current trends and state-of-the-art in RBTG.
3.3. Data Extraction
The 267 selected studies were thoroughly reviewed and inspected, with data extracted according to the research questions: Table 4 summarizes the extracted data. To minimize the risk of missing information and reduce potential errors, we performed reviews both during the data-extraction process, and after its completion.
Table 4. Data Collection for Research Questions
RQs Type of Data Extracted
RQ1 Core information for each paper (publication year, publisher, type of paper).
RQ2 Type of requirements, descriptions, classifications.
RQ3 Type of techniques, degree of automation, intermediate representations, descriptions.
RQ4 Type of generated output artifacts and test execution environment.
RQ5 Tool name, function, language, types, and descriptions.
RQ6 Evaluation methods, metrics.
RQ7 Application domains, programs in case studies.
RQ8 Details of remaining challenges, potential future work.
4. Answer to RQ1: What are the Publication Patterns of RBTG Studies?
267 studies were identified for critical evaluation regarding RBTG, with the extracted data summarized in Table 4. Core information was extracted for each study, including publication year, publisher, and type of paper. This section provides an overview of the surveyed studies, focusing on their publication trends and venues.
4.1. Publication Trends
Figure 2(b) provides a comprehensive overview of the yearly RBTG publication trends of the 267 primary studies, up to the end of 2024: Figure 2(a) shows the number of publications each year; and Figure 2(b) presents the cumulative number of publications over time. The earliest reviewed publication was from 1994. This field has attracted significant attention and interest from researchers since 2001. Since 2008, there have been at least ten published studies each year, with notable peaks of over 20 papers in 2019 and 2021. Figure 2(b) shows the polynomial trendline, with a high determination coefficient ($R^{2}=0.9976$): There is both a sustained and growing interest in RBTG, highlighting its importance in software-testing automation.
(a) Number of publication per year.
(b) Cumulative number of publications per year.
Figure 2. Overview of the surveyed 267 RBTG studies’ publication.
[…]
Overview of the surveyed 267 RBTG studies’ publication up to 2024.
4.2. Distribution of Publication Venues
Figure 3(b) shows the distribution of publications across conferences and journals, with Figure 3(a) presenting the overall venue distribution, and Figure 3(b) illustrating the distribution per year. As shown in Figure 3(a), 175 (66%) of the 267 primary studies were published in conferences and 92 (23%) were in journals. The number of studies published in conferences has consistently exceeded those in journals over the years, reflecting conferences as key platforms for disseminating new research findings quickly and effectively.
(a) Overall venue distribution.
(b) Venue distribution per year.
Figure 3. Venue distribution for RBTG studies.
[…]
5. Answer to RQ2: What Types of Requirements Have Been Used in RBTG?
Requirements play a crucial role in RBTG, as their content and representations directly influence generation strategies and approaches, ultimately affecting the quality of the resulting test cases. In real-world projects, requirements are often written in NL, which is easy to understand, but inherently ambiguous. To address this, NL requirements are often transformed into formal or semi-formal specifications. This section categorizes the collected papers based on requirement types, identifying the most commonly used representations, and analyzing their evolution over time.
As explained in Section 2.1, this paper uses five requirements categories: NL; semi-formal; model-based; formal mathematical; and hybrid. Figure 4(a) presents the distribution of the types of requirements, showing that 53% (141 papers) focused on model-based representations; 28% (75 papers) involved NL; and 19% (51 papers) used other representations. Figure 4(b) shows the requirements’ types over time: Although model-based requirements have been extensively studied over the years, NL types began to attract attention in 2011, and have surpassed model-based approaches since 2019.
(a) Overall requirements types.
(b) Requirements types per year.
Figure 4. Types of requirements.
[…]
Types of requirement.
5.1. Natural Language Requirements
Users usually communicate the needs of a software application in NL. This has meant that generating test cases directly from NL requirements has become a key objective for researchers. Among the surveyed studies, the NL-based requirements can be categorized into three types: unstructured; templated; and controlled natural language (CNL) requirements. Unstructured NL refers to free-text expressions with no constraints on the requirements’ formulation: This is the most commonly used approach in practice. Templated requirements include use case descriptions, user stories (which are commonly employed in agile development), and scenarios. A use case is a sequence of interactions between a system and a user (understood as an instance of an actor) (Ryser and Glinz, 1999). A scenario depicts a portion of the application’s behavior at a specific time and place, capturing the situational context (Sarmiento et al., 2014a): Use cases can be seen as distinct models of scenarios. CNL requirements use restricted grammar and predefined vocabulary to reduce textual complexity and ambiguity.
Figure 5. Distribution of types of natural language requirements.
[…]
Distribution of types of natural language requirements.
Figure 5 presents the distribution of different NL types used in the surveyed studies. Unstructured natural language (Masuda et al., 2021; Arora et al., 2024) represents the main form of input requirements, used by 27 studies (36%). The use case description (Mahalakshmi et al., 2018; Sneed, 2018) is also widely adopted, appearing in 24 studies (32%). Other notable NL types include user stories (seven papers), scenarios (four papers), CNL like SysReq (six papers), and ITLingo RSL (three papers). SysReq is a CNL for system requirements, defined by a phrase-structure context-free grammar and a lexicon encompassing the application domain vocabulary (Carvalho et al., 2014b; Silva et al., 2015). ITLingo RSL (Requirements Specification Language) is a controlled and integrated natural language, designed to assist in the systematic, rigorous, and consistent production of requirements specifications (Paiva et al., 2019; Miranda et al., 2020). Although NLP and ML techniques have also been combined to directly transform these requirements into test cases (Arora et al., 2024; Ueda et al., 2024), the majority of studies use an intermediary step, converting the requirements into structured models before generating the test cases. Further details about these approaches will be explored in Section 6.
5.2. Semi-formal Requirements
Semi-formal requirements combine NL with structured templates, tables, or simplified graphical representations. Examples include ontology-based specifications (Banerjee et al., 2021), SysML (Chen et al., 2023), and courteous logic representation (Sharma and Biswas, 2014b).
Within this category, the requirements ontology (five papers) and SysML (three papers) were the most frequently used. The requirements ontology represents concepts and relationships within a domain (Banerjee et al., 2021; Ul Haq and Qamar, 2019), with inference rules describing strategies for deriving test cases from the ontology (Tarasov et al., 2017). SysML models the system under test (SUT) along with its interactive external environment (Zhu et al., 2021), providing diagrams that enable the textual depiction of requirements, including non-functional ones (Chen et al., 2023; Abbors et al., 2009). Semantic specifications (Moitra et al., 2019; Crapo and Moitra, 2019) that are formulated in the controlled English semantic application design language (SADL) map directly into the web ontology language (OWL), ensuring readability and comprehension for both humans and machines. Courteous logic-based requirements representations (Sharma and Biswas, 2015, 2014a), expressive decision tables (EDTs) (Venkatesh et al., 2015a, b) and domain specific language (DSL) (Cabral and Tamai, 2008; Olajubu et al., 2017) have also been adopted, each appearing in two studies. Despite their potential to streamline test generation, only 22 studies (8%) used semi-formal requirements: These were mostly applied to embedded systems (six papers) and safety-critical domains, like avionics and automotive (seven papers).
5.3. Model-based Requirements
Requirements models provide an intuitive and visual representation of system requirements. This section examines the models used in the surveyed studies. Figure 6 presents the distribution of requirements models across the studies. The majority of papers (72%) used UML diagrams, highlighting the significant focus on UML in the literature. Automata models, such as finite state machines (FSMs) and extended finite state machines (EFSMs), were used in 8% of the papers. Other models included: use case maps (3%), meta-models (3%), business process models (2%), specifications (2%), tree models (2%), B models (1%), and event-B models (1%). There were also several other uncommon models, each appearing only once.
Figure 6. Distribution of model-based types.
[…]
Distribution of model-based types.
5.3.1. UML Diagrams
The UML has been widely used as a system-level specification language. It is a visual modeling language that comprises various diagrams, each representing different aspects of a system’s structure and behavior. Figure 6 shows that UML ADs (Linzhang et al., 2004; Mingsong et al., 2006; Fan et al., 2009) (23%) are the most commonly used model: As a behavioral model, ADs provide detailed implementation information, making them well-suited for test derivation. However, their effectiveness decreases when dealing with large and complex systems. Combinational UML diagrams (14%), sequence diagrams (11%), statechart diagrams (11%) and use case diagrams (11%) have also been widely used. Combinational UML diagrams combine the strengths of multiple models, enabling the generation of more reliable test cases — however, they also incur a time overhead (Pérez and Marín, 2018). Examples of combinational UML diagrams include activity and statechart diagrams (Sahoo et al., 2021; Swain et al., 2010b), sequence and use case diagrams (Sarma and Mall, 2007; Swain et al., 2010a) and sequence and statechart diagrams (Khurana and Chillar, 2015; Sokenou, 2006). Both sequence (Khandai et al., 2011a; Samuel and Joseph, 2008) and statechart diagrams (Hametner et al., 2011; Pradhan et al., 2022) are also behavioral models, capturing system interactions and state transitions to support test generation. In contrast, use case diagrams (Hamza and Hammad, 2021; Nebut et al., 2006) describe system functionality from a user’s perspective, providing a high-level view of interactions without detailing execution flow or state changes. Communication diagrams (Samuel et al., 2007) convey both the messages exchanged between objects and their execution order, providing insight into both data flow and control flow. Similarly, interaction diagrams (Nayak and Samanta, 2009) visually depict the sequence of messages exchanged among collaborating objects across different scenarios of a use case.
5.3.2. Automata
Automata (Hopcroft et al., 2001; Linz and Rodger, 2022) are abstract mathematical machines used to model the behavior of systems, usually through states and transitions between them. Four types of automata were used in the surveyed papers: EFSM (Elqortobi et al., 2023; Rao et al., 2016); FSM (Rechtberger et al., 2022; Liu and Miao, 2010); safety SysML state machine (S2MSM) (Wang et al., 2022b); and Mealy machines (Pfaller, 2008). EFSM was the most commonly used, appearing in five papers, followed by FSM, which appeared in four. Both S2MSM and Mealy machines were found in one paper each. Table 5 compares the different automata models.
Table 5. Automata Models
Automata Model Tuple Structure Mathematical Definition Description
Finite state machine (FSM) 5-tuple $M=(S,s_{0},I,O,T)$ $S$ is nonempty finite set of states, $s_{0}\in S$ is the initial state, $I$ and $O$ denote nonempty finite sets of inputs or outputs respectively, $T$ is the finite set of transitions
Extended finite state machine (EFSM) 6-tuple $M=(S,s_{0},V,I,O,T)$ $V$ is the finite set of (internal) context variables
Safety sysML state machine (S2MSM) 7-tuple $M=(S,T,E,G,P,A,V)$ $E$ is a finite set of events, $G$ is a finite set of guards, $P$ is a finite set of the priority, $A$ is a finite set of actions, $V$ is a finite set of variables
Mealy machine 6-tuple $M=(S,s_{0},I,O,\delta,\lambda)$ $\delta:S\times I\rightarrow S$ and $\lambda:S\times I\rightarrow O$ are total functions which denote the deterministic transition and output relations
An FSM is a finite labeled transition system where each transition is associated with an input/output pair, consisting of an input action and a corresponding output action (Ibias et al., 2021; Liu and Miao, 2010). An EFSM system model is an extension of FSMs that includes additional features, like variables and conditions on transitions (Chen et al., 2018; Vaysburg et al., 2002). S2MSM has additional verification and refinement procedures to check the safety of the generated model using a formal verification tool: S2MSM can both precisely model functional requirements and capture safety concerns, in a unified model (Wang et al., 2022b). A Mealy machine is an FSM where the outputs are determined by both the current state and the current input. It enables compact modeling of reactive systems by associating outputs directly with transitions, allowing for precise representation of input–output behaviors (Pfaller, 2008). Due to their logical structure, these automata are well-suited for modeling the structural and behavioral aspects of software applications, particularly in testing. However, their effectiveness is limited when handling more complex data dependencies and complex systems.
5.3.3. Use Case Map
A use case map (UCM) is a visual notation for modeling scenarios/workflows derived from requirements (Kesserwan et al., 2019). It is similar to a UML AD, but imposes fewer constraints, emphasizing the causal relationships between workflow steps without specifying detailed message exchanges or data (Boucher and Mussbacher, 2017). Four of the surveyed papers used UCMs as input (Boucher and Mussbacher, 2017; Kesserwan et al., 2019; Boulet et al., 2015; Amyot et al., 2005).
5.3.4. Meta-Model
Meta-models standardize the information used in test-case generation: They formally specify a concept by defining its attributes and its relationships with other concepts (Gutiérrez et al., 2012). Gutiérrez et al. (2012, 2015) used a functional requirements meta-model as the basis for testing: The core concept of this meta-model was the functional requirements’ element, defined by (1) step elements representing discrete actions (e.g., data requests or computations); and (2) execution-order elements specifying their sequence. Chen et al. (2022) proposed a test-requirement meta-model that defined the core concepts along with their relationships and dependencies. Granda et al. (2014) applied transformation rules to convert meta-models into test cases.
5.3.5. Other Types of Models
In addition to the models already discussed, several others were also used. Business process models (BPMs) (Yazdani Seqerloo et al., 2019; Khader and Yousef, 2016; Gupta and Surve, 2011) represent a series of related activities producing a specific service or product. A Cause-effect graph (CEG) (Krupalija et al., 2022; Kalaee and Rafe, 2016; Paradkar et al., 1997) is a black-box testing technique that describes relationships between events using logical relations (and events are labeled as either causes or effects). A Behavior tree (BT) model (Lindsay et al., 2015; Salem and Hassan, 2010) is a directed tree with various node types and branching, aiding in modeling complex behaviors. Additional models include: Event-B models (Malik et al., 2009; Satpathy et al., 2006); B models (Bouquet et al., 2005; Legeard and Peureux, 2001); scenario-based state/event trees (Tsai et al., 2003); EDTs (Agrawal et al., 2020); variable relation models (VRMs) (Wang et al., 2020; Zhipeng et al., 2021); and multiple condition-control flow graphs (MCCFGs) (Son et al., 2019). These models all provide different approaches for representing requirements.
5.4. Formal Mathematical Requirements
Formal mathematical specifications use precise logical formulas and constraint-specification languages to express system requirements, typically using mathematical notations (Bruel et al., 2021). A Structured object-oriented formal language (SOFL) (Saiki et al., 2021; Cajica et al., 2021; Sato and Sugihara, 2015) integrates graphical and formal notations, defining functions through pre- and post-conditions. Z formal specification (Kadakolmath and Ramu, 2022; Helke et al., 1997; Chang et al., 1998; Stocks and Carrington, 1996) structures specifications into modular units called schemas using formal mathematical notation. Linear temporal logic (LTL) (Aniculaesei et al., 2019b, a; Whalen et al., 2006) describes the temporal evolution of system behaviors with mathematical formulas: It has been widely used in formal verification, especially for model checking. Other formal specifications, such as software cost reduction (SCR) specifications (Gargantini and Heitmeyer, 1999), functional scenario form (FSF) (Liu and Nakajima, 2020) and featured transition systems (FTSs) (Devroey et al., 2014), have also been used. Although these specifications offer advantages (like automatic generation of test cases), they also face disadvantages in terms of meeting coverage requirements efficiently.
5.5. Hybrid Requirements
The hybrid category includes requirements specifications that combine two or more types of representations. A combined textual specification, for example, may consist of a use case diagram, textual use cases, HTML screen mockups, and a glossary defining key terminology (Clerissi et al., 2017). Other combinations have also been observed, such as UML diagrams integrated with business vocabulary and business rules (Essebaa et al., 2019) or object constraint language (Ali et al., 2014), as well as visual requirement specifications or prototypes (Singi et al., 2015). Among the surveyed papers in this review, seven (3%) used hybrid requirements. Future requirements may likely integrate multiple types of formal specifications (Clerissi et al., 2017): This hybrid approach helps to ensure the correctness and completeness of requirements, supporting the generation of effective test cases.
6. Answer to RQ3: What Are the Approaches that Support RBTG?
RBTG generation approaches take requirements as input and produce test cases as output. Different requirements’ representations use different strategies to generate test cases. Of the 267 surveyed studies, 39% generated test cases directly from the requirements, with the other 61% first transforming the requirements into intermediate representations (before generating test cases). Figure 7 shows that most common forms of requirements (such as models and NL) use intermediate representations as the basis for test-case generation. Accordingly, this section examines these two main strategies: direct test-case generation and test-case generation through intermediate representations.
Figure 7. Test-case generation approaches with/without intermediate representations.
[…]
Test-case generation approaches with/without intermediate representations.
6.1. Test-Case Generation without Intermediate Representations
Table 6. Approaches without Intermediate Representations
Requirements Approach No. of Studies
Natural language Natural language processing (NLP) 7
Machine learning (ML) 2
Model-based approach 2
Hybrid approach 2
Other specific approach 6
Semi-formal Ontology‑based approach 4
Model-based approach 3
Semantic-head-driven approach 2
ASSERT™ 2
Row-guided random algorithm with fuzzing (RGRaF) 2
Model-based Model-based approach 31
Heuristics-based approach 5
Transformation-based approach 3
Hybrid approach 2
Other specific approach 15
Formal mathematical Model-based approach 6
Heuristic-based approach 3
Specific approach 5
Hybrid Other specific approach 2
Table 6 summarizes the direct generation approaches identified in the surveyed papers. These approaches are categorized based on the types of input requirements: The form of the requirements naturally shapes the generation strategy. Within each category, the approaches are arranged in the order of their frequency of occurrence in the reviewed studies.
In the context of “Natural Language” requirements, NLP remains the most widely adopted: This includes techniques such as syntax testing (Intana et al., 2020), which analyzes and captures the sequence of patterns or syntax of variables from the data dictionary; syntactic similarity (Masuda et al., 2016), which determines the case conditions and actions; and keyword extraction (Ansari et al., 2017). Recently, LLMs have gained significant attention and are increasingly becoming a preferred approach for test-case generation (Arora et al., 2024; Mathur et al., 2023; Yin et al., 2024; Shakthi et al., 2024). ML has also been commonly used: Ueda et al. (Ueda and Tsukada, 2021), for example, enhanced the accuracy of automatic test-case generation by training data selection. Similarly, Kikuma et al. (Kikuma et al., 2018) trainind ML on tagged data, and then applied the model to new requirements specifications to generate test cases. Model-based approaches have also been used to derive test cases from CNL requirements, making use of the characteristic CNL constructs. However, these approaches may require substantial manual effort (Paiva et al., 2020; Maciel et al., 2019). Hybrid approaches integrate multiple methodologies: Fernandes et al. (Fernandes and Maciel, 2020), for example, integrated equivalence class partitioning and boundary value analysis, while Elghondakly et al. (Elghondakly et al., 2015) combined text mining with symbolic execution methodologies. Specific tailored approaches have also been proposed: Dwarakanath et al. (Dwarakanath and Sengupta, 2012), for example, introduced Litmus, which uses a syntactic parser (called the Link Grammar parser) alongside pattern-matching rules. These rules were designed to evaluate testability, identify test intents, and systematically generate both positive and negative test cases — a positive test case verifies the system’s expected behavior under normal conditions, and a negative test case checks the system’s response to exceptional or invalid inputs.
“Semi-formal” approaches primarily depend on the structure and form of the input requirements. These approaches include those based on ontologies, such as the approaches by Ul Haq and Qamar (2019) and Feldmann et al. (2014), which integrate a knowledge-based system (ontology) with a learning-based testing algorithm to automate the test-case generation. Similarly, Tarasov et al. (2017) used inference rules to query the ontology, enabling the verification of conditions and the retrieval of the necessary data for test-case construction. Model-based approaches are also included in this category, where test cases are derived from SysML models (Zhu et al., 2021; Abbors et al., 2009). The semantic-head-driven generation approach (Sharma and Biswas, 2015), produced functional test cases using knowledge stored in a courteous logic (Lafi et al., 2021) representation of requirements — this is an expressive subclass of standard logical representations that incorporates procedural attachments for prioritized conflict resolution. Hybrid approaches have also been proposed, such as the $ASSERT^{TM}$ (Moitra et al., 2019; Crapo and Moitra, 2019; Li et al., 2019): This approach uses a suite of tools designed to capture requirements; provides explainable and automated formal analyses to detect and resolve errors during the requirements authoring process; and automatically generates a complete set of test cases. The RGRaF (Venkatesh et al., 2015a, b) applied a dependency-driven random algorithm with fuzzing to generate test cases, targeting time-boundary conditions.
In the “Model-based” category, naturally, model-based approaches represent the main strategy, with the core methodology centered on constructing detailed models derived from requirements. Test cases are extracted using traversal algorithms, such as depth first search (DFS) (Linzhang et al., 2004), pre-order traversal algorithms (Yonathan et al., 2024), and category partition methods (Gutiérrez et al., 2015). Several heuristics-based approaches have also been applied for model-based requirements: Genetic algorithms (Ibias et al., 2021; Shirole and Kumar, 2010), ant-colony algorithms (Agrawal et al., 2020), and metaheuristic techniques (Khalifa et al., 2019) have all been used to generate test cases. A genetic algorithm uses iterative stochastic optimization through selection, crossover, and mutation to evolve test-case populations, converting objectives into quantifiable fitness functions, and evaluating generated test data. An ant colony algorithm, inspired by the way ants use pheromones to find routes between colonies and food, can solve computational problems and identify optimal paths within graphs. Metaheuristics, as problem-independent optimization strategies, are black-box techniques capable of finding global optima without relying on greedy search methods. Transformation-Based Approaches (Boucher and Mussbacher, 2017; Hue et al., 2018) represent specific implementations of model-based strategies, focusing on unique traversal mechanisms. Hybrid approaches combine multiple methods, such as equivalence and classification tree methods (CCTM) (Intana and Sriraksa, 2019), to improve testing efficiency and coverage.
Model-based and heuristic-based techniques are common in “formal mathematical” approaches. Kadakolmath and Ramu (2022) introduced a ProZ model-checking-based approach that uses a breadth-first search strategy to generate test cases. Aniculaesei et al. (2019b) used the NuSMV model checker to derive test cases from LTL requirements. Liu and Nakajima (2020) proposed a vibration method to heuristically generate test cases, and Wang et al. (2019) used a genetic algorithm to identify the best mutant test condition, which was then used to generate test data.
Among the approaches that generate test cases directly from requirements (without intermediate models), model-based approaches are very versatile and applicable across multiple requirements’ forms. They are also the most widely adopted strategy. In addition to model-based techniques, heuristic-based approaches and NLP are also frequently used. Among the heuristic-based approaches, genetic algorithms are particularly prominent. Recently, NLP has become an increasing popular technique for test-case generation.
6.2. Test Case Generation with Intermediate Representations
Many test-case generation approaches use an intermediate representation as part of the process: This helps with the extraction and formalization of key information from requirements, improving both the accuracy and effectiveness. As Figure 7 shows, 61% of the surveyed studies use an intermediate model as part of the transformation process.
6.2.1. Intermediate Representations
This section explores the different types of intermediate representations commonly used in the reviewed studies.
Table 7. Intermediate Representations
Intermediate Representation
Categories
Intermediate Representations No. of Studies Total
Graphical models Activity diagram (AD) or activity graph (AG) 19 90
Combined directed graphs 10
Control flow graph (CCFG) 9
Sequence diagram (SD) or sequence graph (SG) 8
Trees 7
State shart diagram or statechart diagram graph (SCDG) 6
State transition diagram (STD) 4
Cause-effect-graph (CEG) 3
Use case test models (UCTMs) 3
Flow dependency graph (FDG) 2
Other diagrams or other directed graph 19
State/Process models Finite state machine (FSM) 11 47
Communicating sequential processes (CSP) test models 6
Extended finite state machine model (EFSM) 4
Meta-model 2
Symbolic model 2
Knowledge graph 2
Other specific models 20
Logical specifications Functional scenarios form (FSF) 3 20
Requirement interfaces 2
Requirement specification ontology 2
Software cost reduction (SCR) specifications 2
Other formalization 11
Petri nets Coloured petri nets (CPN) 2 4
Petri-net models 2
Tabular formats Activity dependency table (ADT) 1 3
Decision table 2
Table 7 groups the intermediate formats into five categories: graphical models; state/process models; logical specifications; Petri nets; and tabular representations. Graphical models include diagrams and directed graphs that represent system states and transitions. State/process models are formal models that describe system behavior with mathematical notations. Logical specifications use specification-based expressions to define requirements. Petri nets, with their unique notation and formal semantics, are treated as a separate category. Finally, tabular representations arrange requirements’ logic into structured tables for enhanced clarity and interpretability. As Table 7 shows, graphical models were the most frequently utilized intermediate format, followed by state/process models. Logical specifications, Petri nets, and tabular formats are comparatively less common.
ADs (Sarmiento et al., 2014a; Hamza and Hammad, 2021; Sarmiento et al., 2014b) or activity graphs (Chen et al., 2023; Kamonsantiroj et al., 2019; Swain et al., 2013) are the most commonly used graphical model: This aligns with the findings in Section 5.3. The second most popular models were combined directed graphs (10 papers): These integrate multiple representations, such as control-flow graphs (CFG) and data-flow graphs (DFG) (Elqortobi et al., 2023; El Qortobi et al., 2020), or statechart diagram graphs and sequence-diagram graphs (Lakshminarayana and SureshKumar, 2020). The third most popular were control flow graphs (nine papers) (Lafi et al., 2021; Swain et al., 2010b; Bansal and Sabharwal, 2013), which provide a static representation of models by capturing all possible control flow paths. Other graphical formats include sequence diagram/graph (Ekici and Tuglular, 2023; Panigrahi et al., 2018; Dhineshkumar et al., 2014), trees (Gupta et al., 2023; Cunning and Rozenbiit, 1999; Tiwari and Gupta, 2013), SCDG (Pradhan et al., 2022; Fröhlich and Link, 2000), STD (Chatterjee and Johari, 2010), CEG (Fischbach et al., 2023), UCTMs (Wang et al., 2022a), and FDG (Kaur and Vig, 2018). Additionally, 19 other studies used various other specialized diagrams or directed graphs tailored to specific contexts (Wang et al., 2022b; Tiwari and Gupta, 2015; Yin et al., 2017; Khurana et al., 2016; Li et al., 2013).
FSMs were the most widely adopted state/process models, with 11 studies exploring their use (Ali et al., 2021; Clerissi et al., 2017; Bin and Anbao, 2012). Communicating sequential processes (CSP) models (Nogueira et al., 2019, 2016, 2014) were also frequently used for formalizing system behaviors. Other intermediate models include EFSM (Rocha et al., 2021; Jiang and Ding, 2011), meta-models (Allala et al., 2022, 2019), symbolic models (Schwarzl and Peischl, 2010), and knowledge graphs (Nayak et al., 2020; Verma and Beg, 2013). Various other specialized models were also explored (Fujita and Ueda, 2024; Bhatt et al., 2022).
Other intermediate representations that were less common included logical specifications (Saiki et al., 2021; Sato, 2020; Carvalho et al., 2014b; Chinnaswamy et al., 2024; Charoenreh and Intana, 2019; Carvalho et al., 2013), Petri nets (Silva et al., 2019; Sarmiento et al., 2016; Reza and Kerlin, 2011), and tabular representations (Noikajana and Suwannasart, 2008). However, even though less frequently used, these formats offer unique advantages, and may support specialized test-generation approaches.
In summary, the transformation of raw requirements into intermediate representations has often employed specialized algorithms. Diagrams and directed graphs are the most popular choice, highlighting their effectiveness in test-case generation. However, alternative intermediate formats (such as formal models and formal specifications) may advance test-generation methodologies and enhance their applicability.
6.2.2. Test Case Generation from Intermediate Representations
Methods for generating test cases based on intermediate models align closely with the “model-based” category in Table 6. Model-based approaches (Khandai et al., 2011b; Nayak and Samanta, 2011) dominate this field, followed by heuristics-based methods (Shirole et al., 2011). These studies highlight that, in most research, requirements are first transformed into specific models, which then support the efficient extraction of the test cases: Various algorithms and techniques are then applied to generate the test cases, such as traversal algorithms (Nogueira et al., 2014; Dalai et al., 2012; Achimugu et al., 2021; Septian et al., 2017), category-partition method (Swain and Mohapatra, 2010; Hartmann et al., 2000; Hierons, 1997), vibration method (Saiki et al., 2021; Liu and Nakajima, 2011), and other specialized techniques (Almeida et al., 2013; Pechtanun and Kansomkeat, 2012; de Figueiredo et al., 2006).
7. Answer to RQ4: What Types of Test-Case Representations Have Been Explored in RBTG?
Test cases can be generated using different approaches and representations, which then structure the outputs into various forms. This section examines on the types of test cases produced. Through an analysis of the reviewed papers, we identified three primary categories: Abstract test cases (ATCs); Concrete test cases (CTCs); and Test scenarios (TSs). (Detailed definitions of these categories were provided in Section 2.2.)
The distribution of each type of test output is shown in Figure 8(a). ATCs make up the majority of outputs, accounting for 61% of the reviewed studies. CTCs make up 33%, and can be executed either manually or automatically. Finally, TSs represent only 6% of the test outputs. In this paper, studies that specified the manual conversion of ATCs into CTCs were categorized as ATCs rather than CTCs: This helps to distinguish between test outputs that are directly executable and those requiring intermediate manual processing. As shown in Figure 8(b), ATCs account for more than 60% of the generated outputs for NL, semi-formal, and model-based requirements; ATCs represent 57% of outputs from hybrid requirements. Requirements expressed through formal mathematics, in contrast, result in less than 32% ATCs: This is because mathematical notations and applications in fields such as signal processing or control systems often define precise input-output relationships, making them more suitable for the direct generation of CTCs, rather than ATCs.
(a) Overall distribution of test-case types.
(b) Test-case distribution by requirement input.
Figure 8. Distribution of different test types.
[…]
Distribution of generated test cases.
ATCs are high-level descriptions of testing requirements that are human-readable. They cannot be directly executed: They require additional modifications, such as incorporating concrete test data and defining test oracles, to transform them into executable test cases. The reviewed papers identified several common representation types for ATCs, including: textual test cases, test paths, tabular test cases, test cases in other formats.
  • Textual test cases: These closely resemble manually written test cases, and are easy for humans to understand and analyze. Textual ATCs are often derived from NL or semi-formal input requirements. Yin et al. (2024) used LLMs to generate textual ATCs from NL requirements. Fujita and Ueda (2024), Ueda and Tsukada (2021), and Fujita and Ueda (2024) used ML techniques to generate textual test cases. Crapo and Moitra (2019) and Sharma and Biswas (2015) proposed approaches to obtain test cases from semi-formal requirements.
  • Test paths: Test paths are derived by traversing specific graphs or models using specialized algorithms. Gupta et al. (2023), for example, generated test paths from root to leaf nodes. Hamza and Hammad (2019), Kamonsantiroj et al. (2019) and Kim et al. (2007) extracted test paths from activity graphs. Lafi et al. (2021) and Alrawashed et al. (2019) derived test paths from control flow graphs.
  • Tabular test cases: These use decision tables with test inputs and outputs being represented as true/false values. Aoyama et al. (2020) and Silva et al. (2019) derived test cases by defining input steps based on the true or false conditions specified in the decision tables. Almohammad et al. (2017), Olajubu et al. (2017), and Kalaee and Rafe (2016) also used decision tables for test-case representation. The tabular form is particularly useful for industrial control and embedded systems, where inputs and outputs are often digital signals.
  • Test cases in other formats: There are also several other less common representations. Gnesi et al. (2004), for example, generated test cases in a language combining elements of process algebra that offered a formal description of system behavior. Chinnaswamy et al. (2024) used regular expressions, and Kadakolmath and Ramu (2022) constructed test cases using specified predicates, storing them in XML format.
CTCs are low-level test cases with concrete (implementation-level) values as input data and expected results. CTCs are more closely aligned than ATCs with actual execution. They can be executed (manually, or automatically using test scripts) in a specific environment.
  • Manual execution: Manually-executed CTCs often augment ATCs with detailed test data. The data can be generated through manual input, randomization, or specialized methods tailored to specific applications. Gupta and Mahapatra (2023) and Wang et al. (2015) used semi-automatic techniques to generate textual CTCs that were then manually executed. Sun et al. (2016), Jena et al. (2014), and Zhang et al. (2014) developed test-case generation approaches by first randomly filling values along test paths without considering input-variable constraints, and then selecting cases that met the guard conditions.
  • Automated execution: Automated CTC execution is primarily driven by test scripts, which can be generated using advanced tools and frameworks. Fernandes and Maciel (2020) used AutomTest, a tool developed in Python 3, to produce executable CTCs. Aoyama et al. (2021) and Khan et al. (2022) converted ATCs into PROMELA test programs for their specific test execution environments. Granda et al. (2021) generated test scripts based on relations between requirements, GUI elements, and Sikulix code in the Eclipse editor. Maciel et al. (2019), Clerissi et al. (2017), and Amalia et al. (2022) used Selenium WebDriver to create executable test scripts, and Pradhan et al. (2022) implemented test cases within the TestNG framework of J-Unit.
TSs are documents specifying a sequence of actions for the execution of a test. Arora et al. (2024) noted that test scenarios simulate diverse user interactions and system behaviors. These scenarios can be represented in various formats, such as step sequences in UML activity diagrams, sequence diagrams, or NL descriptions. Some and Cheng (2008) generated test scenarios based on path sequences, each comprising a test-scenario identifier, a setup section, a sequence of test steps, and an expected outcome.
In summary, test cases can be derived as ATCs in textual form or as textual test paths generated through graph traversals. However, these test cases often lack sufficient detail and completeness, and are unready for direct execution. CTCs can be generated semi-automatically with human intervention, bridging some of these gaps. Nevertheless, achieving a fully automated pipeline for both test-case generation and execution remains essential to significantly enhance testing efficiency and minimize manual effort. This transition from ATCs to CTCs poses substantial challenges and is an area of ongoing research.
8. Answer to RQ5: What Tools Have Been Developed to Support RBTG?
Across the reviewed literature, we identified 88 tools designed specifically to support RBTG tasks, collectively referred to as RBTG tools. Table 8 classifies these RBTG tools into two groups, according to their operational phases: requirements processing, and test generation. (If a tool supports both requirements processing and test-case generation, it was classified under the test generation category.) The table includes the tools’ names (e.g., jUCMNav), relevant reference (e.g., [6]), and implementation language (e.g., Java): Interested readers can find the original studies by consulting the references.
Table 8. RBTG Tools
Type Tools Counts
Requirements Processing
Open Source SequenceDiagram2ESG (Ekici and Tuglular, 2023) (Java) , spaCy (Gröpler et al., 2021) (Python), ASSERT™ tool chain (Moitra et al., 2019) (Not specified), Text2Test (Bhatt et al., 2022) (Not specified), USLTG (Hue et al., 2019) (Java), jUCMNav (Boulet et al., 2015; Kesserwan et al., 2023) (Java), SystemCockpit (Aichernig et al., 2014) (Not specified), MoMuT::REQs (Aichernig et al., 2014) (Not specified), ROO (Intana et al., 2023) (Java), ArgoUML (Sun et al., 2009) (Not specified), AToM3 (Hettab et al., 2015, 2013) (Python), Sahi (Boghdady et al., 2011a) (Not specified), Semiformalizer (Aoyama et al., 2020) (Not specified) 13
Research Prototype GUI-Test tool (Granda et al., 2021) (Java), DODT (Aichernig et al., 2014) (Not specified), RETNA (Boddu et al., 2004) (Java), XML-based tool for State/Event Tree (Tsai et al., 2003) (Java), TestMEReq (Moketar et al., 2016) (Not specified), CNLParser (Carvalho et al., 2014a) (Not specified), IMR-Generato (Carvalho et al., 2014a) (Not specified), CAT (Lu, 1994) (Not specified), VDM (Dick and Faivre, 1993) (Not specified) 9
Commercial IBM Engineering Requirements Management DOORS (Moitra et al., 2019) (Java), IBM Rational Rose (Kaur and Vig, 2018; Kansomkeat and Rivepiboon, 2003) (Java), Magic Draw UML (Singh and Sharma, 2015; Sarma et al., 2007) (Java) 3
Test Generation
Open Source UMTG toolset (Wang et al., 2022a) (Java), Test Generator for Cause Effect Graphs (TOUCH) (Krupalija et al., 2022) (Java), TESTSD2EFSM (Rocha et al., 2021) (Java), AutomTest (Fernandes and Maciel, 2020) (Java), TCGen (Pérez and Marín, 2018) (Java), SPECMATE (Fischbach et al., 2020) (Java), Test-o-Matic (Masud et al., 2017) (Java), REBATE (REquirements Based Automatic Testing Engine) (Sinha et al., 2016) (Java), AndroMDA (Elallaoui et al., 2016) (Java), FDR (Nogueira et al., 2019) (C++), TaRGeT (Nogueira et al., 2019) (Java), Use Case Acceptance Tester (UCAT) (El-Attar and Miller, 2010) (Java), SMT solver Z3 (Aichernig et al., 2015) (C++), EclEmma (Malik et al., 2009) (Java), T-VEC tool (Carvalho et al., 2014b) (Not specified), TestGen (Gutierrez et al., 2008) (Java), JWebUnit (Gutierrez et al., 2008) (Java) 17
Research Prototype MC/DC-TG-RT (Aoyama et al., 2020) (Java), ReqOntoTesGen (Intana et al., 2023) (Java), Tsgenerator (Yang et al., 2021) (Java), SYNTest (Intana et al., 2020) (JavaScript), REQCAP (Daniel et al., 2018) (Not specified), SAGA (Daniel et al., 2018) (Not specified), Virtual Test Engineer (VTE) (Anbunathan and Basu, 2016) (Java), ConcurTester (Sun et al., 2016) (Java), ACT (Abstract to Concrete Tests) (Bubna and Chakrabarti, 2016) (Java) , MISTA (Mirza and Khan, 2018) (Java), UBTCG (Oluwagbemi and Asmuni, 2015) (Java), NAT2TEST (Carvalho et al., 2015) (Java), BT Analyser (Lindsay et al., 2015) (Not specified), C&L (Scenarios & Lexicons) (Sarmiento et al., 2014a) (Lua language), Litmus (Dwarakanath and Sengupta, 2012) (.NET), SOLIMVA (de Santiago Júnior and da Silva, 2017) (Not specified), STATEST (Chatterjee and Johari, 2010) (Not specified), ATCGT (Chen and Li, 2010) (Not specified), Comprehensive Test (ComTest) (Swain et al., 2010b) (Java), UTCG (Swain et al., 2010a) (Java), TSGen (Sun et al., 2009) (Java), ASSIST (Sarma and Mall, 2009) (Java), GTSC (Santiago Júnior and Vijaykumar, 2012) (Not specified), TAD (Noikajana and Suwannasart, 2008) (C# ), TDE/UML (Hasling et al., 2008) (Not specified), UML behavioural test case generator (UTG) (Samuel and Mall, 2009; Samuel et al., 2008; Samuel and Mall, 2008) (Java), GenTCase (Ibrahim et al., 2007) (C++), Condado (Santiago et al., 2006) (Not specified), UC-SCSystem (Nebut et al., 2006) (Java), TSGAD (Xu et al., 2005) (Not specified), LEIRIOS Test Generator (LTG) (Bouquet et al., 2005) (Not specified), ScenTED-DTCD tool (Reuys et al., 2005) (Not specified), UMLTGF (Linzhang et al., 2004) (Not specified), AGATHA (Lugato et al., 2004) (Not specified), EXTRACT (Yu et al., 2003) (Not specified), UCTSystem (Nebut et al., 2003b) (Java), UMLTest (Offutt and Abdurazik, 1999) (Not specified), Objecteering CASE (Nebut et al., 2003a) (Not specified), PerformCharts (Santiago et al., 2006) (Not specified), UsageTester (Riebisch et al., 2002) (Not specified), TCaseUML (Sun, 2008) (Not specified), Requirements Model Code Synthesis (RMCS) (Gupta et al., 2001) (Not specified), RT-Tester (Carvalho et al., 2014a) (Not specified) 43
Commercial Spec explorer (Yazdani Seqerloo et al., 2019) (C#), Conformiq Creator (Hussain et al., 2018) (Java), Conformiq’s Qtronic (Abbors et al., 2009) (Java) 3
(a) Tools types.
(b) Tools’ development language.
Figure 9. Distribution of RBTG tools
[…]
Distribution of RBTG Tools.
In the requirements-processing category, open source tools are ranked highly (13 tools), followed by research-prototype tools (nine ) and commercial tools (three). In the test-generation category, research-prototype tools lead significantly (43), with open-source (17) and commercial (three) tools being less common.
Overall, as shown in Figure 9(a), research-prototype tools were the most common RBTG tools, followed by the open source, and then commercial tools. Regarding the programming languages used in the implementation of these tools, as shown in Figure 9(b), Java stands out as the most widely used, accounting for 42 tools. 36 tools did not have their implementation languge specified. The other languages (C++, C#, Python, Lua, .NET, and JavaScript), were used in significantly fewer cases.
9. Answer to RQ6: How Have RBTG Approaches Been Evaluated?
9.1. Overall Evaluation in RBTG
Chen and Babar (2011) described nine types of evaluation methods for software product lines. Four of these were analyzed statistically based on the papers reviewed in this study. To ensure clarity, we renamed the evaluation types as Example, Empirical Case Study, Real-World Case Study, and Discussion.
  • Example. In the example category, an example is used to demonstrate how the proposed method works. The example is primarily illustrative and not intended to validate or evaluate the approach. For instance, Ueda et al. (2024) provided an example covering the entire process from requirements specification through data processing to test-case generation and evaluation. Yonathan et al. (2024) used the login feature as an example to illustrate both the process and the results. Hong et al. (2001) did the same using a coffee-vending machine scenario.
  • Empirical Case Study. This explores a phenomenon within its real-life context. While based on real scenarios, these studies are typically conducted in controlled academic settings, and explicitly designed as case studies. Gupta et al. (2023) demonstrated their technique using a case study of ATM cash-withdrawal processes.
  • Real-World Case Study. In this category, approaches are applied to real-world systems, often in collaboration with companies or public organizations. These studies emphasize practical applicability, and evidence of success in operational environments. Arora et al. (2024) collaborated with Austrian Post, an international logistics and service provider, to generate test scenarios from NL requirements. Intana et al. (2023) demonstrated the effectiveness and accuracy of their approach and tool through two case studies involving real-world systems.
  • Discussion. Discussions provide some qualitative, textual, and opinion-based evaluation, without concrete examples or empirical data. Elqortobi et al. (2023) and Lafi et al. (2021), for example, analyzed the test-generation process primarily through discussions, without providing illustrative examples.
(a) Empirical evaluations.
(b) Evaluation methods.
Figure 10. Overall distribution of evaluation methods.
[…]
The 267 studies were classified based on their primary evaluation methods. Figure 10(a) shows that examples and empirical case studies were the largest proportion, each accounting for 36% of the total. In contrast, only 17% of the studies used real-world case studies (Schnelte, 2009), and 11% were categorized as discussions.
We categorized the reviewed studies according to their validation methods into four distinct types: theoretical demonstration, self-evaluation, human-involvement evaluation, and comparative controlled experiment.
  • Theoretical Demonstration. This category included studies that primarily introduce and illustrate methods through theoretical explanations and examples, without any formal evaluation. Chen et al. (2023), Kaur and Singh (2014), and Reuys et al. (2003) presented their test-generation process without conducting any evaluation.
  • Self-Evaluation. This category included research that performs an initial assessment of the proposed approach using relevant data, typically focusing on internal validation without comparison to other methods. Gupta et al. (2023) and Thanakorncharuwit et al. (2016) evaluated their generated test cases based on some types of test coverage, while Wang et al. (2019) involved mutation testing in the evaluation.
  • Human-Involvement Evaluation. This category involved studies incorporating human participation, either through manual assessment of results, or by comparing automated outputs with expert-generated ones. Arora et al. (2024), Shah et al. (2019), and Fernandez-Sanz and Misra (2012) had their proposed approach critically evaluated by experts. Ekici and Tuglular (2023) compared their automatically generated test sequences with those created manually.
  • Comparative Controlled Experiment. The final category included research that had controlled experiments comparing the proposed method against existing approaches. This category provided rigorous validation of the proposed methods’ effectiveness and advantages. Agrawal et al. (2020) conducted experiments with three other algorithms, using corresponding parameter values for comparison. Elghondakly et al. (2016) and Boghdady et al. (2012a) computed cyclomatic complexity values to measure the number of logical paths in a module and compared these values with those obtained from other methods.
Figure 10(b) shows that theoretical demonstration was the most common validation method, accounting for 52% of the studies. Self-evaluation, human-involvement evaluation, and comparative controlled experiments all had similar representation, with 20%, 16%, and 12% of the studies, respectively. Notably, this suggests that over half of the reviewed studies lack formal evaluation.
In summary, the dominance of examples and empirical case studies indicates that many proposed approaches lack sufficient practical validation: Only 17% of the studies employed case studies based on real-world systems, with most relying on theoretical demonstrations. Only 12% of the studies conducted comparative controlled experiments. The limited use of formal validation methods highlights a significant gap between theoretical research and practical applicability, making it difficult for practitioners to rely solely on existing research for decision making.
9.2. Evaluation Methods and Metrics in RBTG
Various methods have been applied to evaluate proposed RBTG approaches. A range of metrics have been used to measure the efficiency and effectiveness of the generated test cases. This section provides a comprehensive overview of these evaluation methods, including the key metrics used in the reviewed studies.
Figure 11 summarizes the metrics adopted in the surveyed literature, categorized into three groups: Efficiency, Effectiveness, and Other. Similar metrics, such as fault detection, code coverage, and model coverage, were grouped together within the categories. A single paper may have used multiple metrics in the evaluations.
Figure 11. Distribution of evaluation metrics.
[…]
Evaluation metrics.
The effectiveness metrics were categorized into four types: fault detection, model coverage, code coverage, and precision, recall, and F-measure.
  • Fault detection. This metrics include mutation score (Liu and Nakajima, 2020; Xu et al., 2014; Kansomkeat et al., 2010) and detected faults (Ekici and Tuglular, 2023; Kundu and Samanta, 2009). The mutation score, calculated as the ratio of killed mutants to the total number of mutants, is a widely used measure in mutation testing (Liu and Nakajima, 2020). The detected faults metric quantifies the number of defects identified by the generated test cases, providing a direct assessment of their fault-detection capability.
  • Model coverage. This category includes metrics such as transition coverage (Gupta et al., 2023), cyclomatic complexity (Yonathan et al., 2024; Chouhan et al., 2012), activity coverage (Gupta et al., 2023), action coverage (Ali et al., 2021), and fitness value (Lakshminarayana and SureshKumar, 2020). It also includes other coverage criteria, such as state coverage (Pradhan et al., 2022) and key path coverage (Chen et al., 2010). Coverage measurement, which evaluates whether or not the model can reach a certain state or transition, is a commonly used evaluation method.
  • Code coverage. This category includes metrics such as path coverage (Gupta et al., 2023), MC/DC coverage (Wang et al., 2022b; Zheng et al., 2021; Rayadurgam and Heimdahl, 2001), branch coverage (Wang et al., 2022a), and row coverage (Agrawal et al., 2020). It also includes other coverage criteria, like predicate coverage (Swain et al., 2010b). Code coverage measurements have also been frequently used.
  • Precision, recall, F-measure. The F-measure, also known as the F1 Score, combines precision and recall and is used to measure the accuracy of automatic test-case generation (Ueda et al., 2024). The formula for the F-measure is as follows:
    (1) ${Precision}=\frac{{TP}}{{TP}+{FP}}$
    (2) ${Recall}=\frac{{TP}}{{TP}+{FN}}$
    (3) ${F-measure}=\frac{2\times{Precision}\times{Recall}}{{Precision}+{Recall}}$
    (where TP is the True Positive, FP is the False Positive, and FN is the False Negative).
The evaluation of RBTG testing efficiency primarily used two key metrics: time spent to generate test cases (Ali et al., 2021), and number of generated test cases (Chen et al., 2022). These metrics are commonly used to assess the reduction in manual effort needed for effective test-case generation.
Manual comparison metrics (Arora et al., 2024) involve human evaluation based on criteria such as relevance, coverage, correctness, understandability, and feasibility. This can provide qualitative insights into the effectiveness of the generated test cases, complementing quantitative evaluation approaches.
Other studies also adopted specialized evaluation methods, such as comparisons with manual evaluation (cross-validated by two evaluators) (Lim et al., 2024) and proof-of-concept evaluation (Daniel et al., 2018), to assess the feasibility and effectiveness of the proposed approaches.
10. Answer to RQ7: In What Domains and Industrial Applications Has RBTG Been Applied?
The SUT application domain strongly influences RBTG, impacting the formulation and representation of requirements, and the selection of test-generation approaches and tools. Test cases can differ significantly between domains in terms of number, style, detail, and granularity. This section identifies the key application domains that have been explored for RBTG.
228 of the 267 reviewed papers explicitly referenced specific application domains: These were incorporated into the analysis. For papers lacking explicit domain specifications, their case studies’ domains were used. The resulting distribution of application domains is presented in Figure 12. The most extensively studied domain was object-oriented programs, in 63 papers. Other notable domains were embedded systems (46 papers), multi-domains (38 papers), and web applications (26 papers). The term multi-domain refers to approaches applied across multiple system types, such as both object-oriented programs and embedded systems (Boghdady et al., 2012b, 2011b). This highlights the significance of object-oriented programs in RBTG research, with an emphasis on model-based approaches that typically employ UML diagrams to describe object-oriented systems.
Figure 12. Distribution of RBTG application domains.
[…]
Distribution of domains in RBTG.
In object-oriented programs, objects interact to execute specific behaviors (Kaur and Vig, 2018). Within the context of RBTG research, a broad range of applications have been modeled as object-oriented programs, including for: library management (Intana et al., 2023, 2020; Briand and Labiche, 2002), online conference reviewing (Granda et al., 2014), kidney-failure diagnosis (KFD) (Intana et al., 2023), university education (Yazdani Seqerloo et al., 2019), and course registration (Kumaran et al., 2011). This diversity in domains highlights the flexibility and adaptability of object-oriented modeling for RBTG.
Embedded systems, which typically operate with continuous digital input and output signals, have also been extensively explored in RBTG surveys. In such systems, input data are generally collected from sensors as the primary data source, and output data are transmitted to system actuators to perform control actions (Silva et al., 2019). Examples of embedded systems studied in RBTG research include systems for elevator control (Chen et al., 2023), urban railway interlocking (Kadakolmath and Ramu, 2022), BodySense (a safety-critical automotive software) (Wang et al., 2022a), railroad crossings (Krupalija et al., 2022) and medical equipment (Tahat et al., 2001). These examples highlight the key role of embedded systems across safety-critical and control-oriented applications.
Web applications have also been a focus of RBTG research due to their widespread use in real-world systems. Example web applications include: the automatic exam-paper evaluation platform AutoEval (Mathur et al., 2023); the online trading platform Online Broker (Sarmiento et al., 2016); and the hotel-management system Hotel Ambassador (Gutiérrez et al., 2015). These studies highlight the applicability of RBTG to software systems with the kind of dynamic and user-driven functionality typical of web-based applications.
Avionics and automotive domains have also drawn considerable attention in RBTG research, with 14 and 12 papers, respectively. Some examples in these domains include the Avionics Launch Abort System (Wang et al., 2022b), Aircraft Engine Sensor (Moitra et al., 2019) and Adaptive Cruise Control System (Aniculaesei et al., 2018).
Several other application domains have also been explored: Mobile applications, such as the Uber app (Ali et al., 2021), are linked to 10 studies. Numerical programs, including the greatest common divisor (GCD) (Cajica et al., 2021) and Mod process (Wang et al., 2019), were examined in nine papers. Industrial control systems (Crapo and Moitra, 2019) and telecommunications (Kikuma et al., 2018) have been explored in four and three papers, respectively. Other less explored domains, such as healthcare (Löffler et al., 2010), desktop applications (Granda et al., 2021), and logistics (Arora et al., 2024) were each covered in a single paper. These numbers highlight the applicability of RBTG across a wide range of industries and use cases.
Overall, while object-oriented programs, embedded systems, and web applications dominate the RBTG research, the diversity of applications highlights the adaptability of RBTG approaches across different domains. However, the limited exploration of some domains suggests opportunities for further research to enhance test-case generation techniques in these underrepresented areas.
11. Answer to RQ8: What Are the Remaining Challenges and Future Work for RBTG?
Following the analyses in the previous sections, this section summarizes remaining challenges and potential future work for RBTG.
11.1. Challenges
This survey paper has identified that RBTG has attracted considerable research interest and has made notable progress in recent years. However, compared to some fields (like automated code generation), its advances remains limited, and the level of automation achieved is still lacking. Many challenges and open questions remain to be addressed.
11.1.1. Challenge 1: Constructing High-Quality Requirements
Requirements serve as the foundation of RBTG, directly influencing the strategies for test-case generation, and determining the quality of the resulting test cases. In industrial practice, most requirements are expressed in NL, which can face lexical ambiguities, contextual inconsistencies, and structural incompleteness (Mustafa et al., 2021): Test cases thus generated are often unreliable, significantly impacting the overall quality of the software testing and development. In traditional development processes, ensuring the precision, completeness, clarity, and consistency of the requirements often involves iterative reviews, conducted manually by engineers during the early project stages. The requirements can then be continuously refined throughout the development and testing phases to maintain their quality and adaptability.
The reviewed RBTG research revealed that efforts have been made to mitigate NL-related issues by translating into semi-formal (Aoyama et al., 2020) or formal representations, or through mapping tables for NLP (Lafi et al., 2021). However, these approaches often require substantial manual effort (Granda et al., 2021), risk losing important information (such as non-functional requirements), and do not guarantee that the requirements will be complete and unambiguous.
To address these limitations, a potential research direction involves integrating requirements engineering with automated testing to establish high-quality, machine-readable requirements for test-case generation. This could involve the use of prototypes, scenarios, stories, and domain knowledge to construct comprehensive and precise requirements. Advances in NLP and LLMs may also offer potential solutions for refining NL requirements.
11.1.2. Challenge 2: From Abstract to Executable Test Cases
CTCs are executable, and their automatic generation represents the ultimate goal of RBTG. However, the surveyed studies reveal that RBTG primarily produces ATCs. These generated test cases are often incomplete, and lack sufficient detail: This makes them unready for direct execution. To address this, researchers have explored methods to convert ATCs into CTCs.
Wang et al. (2022a) used a mapping table to develop an executable test generator that can automatically produce CTCs. The table maps high-level operation descriptions and test inputs to concrete driver functions and inputs. Maciel et al. (2019) refined ATCs by assigning values to entities and creating temporary variables. These were then transformed into concrete test scripts by establishing relations between the requirements and the syntax of the Robot framework (Gupta and Surve, 2011), using keywords from the Selenium library. Kesserwan et al. (2019) semi-automatically mapped abstract Test Description Language (TDL) datasets to concrete data, generating executable test procedures in TTCN-3 (Kesserwan et al., 2023) for embedded systems.
A common approach for the creation of concrete test data and oracles involves random generation or manual crafting. Fernandes and Maciel (2020) proposed AutomTest, a tool that enables users to generate automated tests from functional requirements: It accepts data within parameter ranges and produces JUnit test cases. However, its applicability is constrained by limited variable support and the need for manual intervention. Pradhan et al. (2022) developed test cases using TestNG: automatically, based on their specific conditions. Moitra et al. (2019) developed executable test scripts tailored for safety-critical systems: These tests can be run using Safety-critical application development environment (SCADE) coverage tools. However, these approaches are often based on formal requirements, and constrained to specific domains.
Exploring a general approach that could automatically convert ATCs to CTCs remains a significant challenge: Addressing these issues will be essential for advancing RBTG.
11.1.3. Challenge 3: Lack of Standard Benchmarks
Research benchmarks enable robust analysis and fair comparisons of automated test-generation approaches. They typically include systems specifically designed for a particular type of evaluation, facilitating comparisons with other research, reducing system-selection bias, and demonstrating the effectiveness of proposed techniques (Fontes and Gay, 2023). Currently, the ATM system (Sahoo et al., 2021) is the most widely-used benchmark. However, the majority of surveyed studies rely only on their own case studies, with nearly half of the reviewed research not involving any case-study validation.
The absence of standardized benchmarks makes it difficult to compare different approaches, reproduce results, and assess real-world applicability. To address this gap, well-documented benchmarking datasets should be developed, incorporating diverse requirements formats (e.g., NL, semi-formal, and model-based). These benchmarks should cover various domains (including embedded systems and enterprise applications) to help ensure consistent evaluation, and support progress in RBTG.
11.1.4. Challenge 4: Rigorous and Efficient Evaluation
Figure 10(b) showed that only 12% of the reviewed studies included controlled comparative experiments, and 16% required human involvement, introducing subjectivity into the evaluation. Current evaluation metrics (such as code coverage, model coverage, and fault-detection rates) may only provide a limited perspective on the effectiveness of RBTG approaches: They may fail to capture how well generated test cases align with the original requirements, especially for non-functional requirements such as performance, security, and usability.
The lack of a comprehensive evaluation framework hinders efforts to determine whether or not generated test cases adequately capture the intended SUT behaviors. To advance RBTG research, rigorous evaluation methodologies are required. These methodologies should extend beyond traditional metrics, incorporating both functional and non-functional dimensions, to enable a more comprehensive assessment of test-case generation techniques.
11.1.5. Challenge 5: Real-World Industrial Implementation
Despite advances in RBTG, its adoption in industrial settings remains limited. As discussed in Section 9, only 17% of the reviewed papers used real-world case studies to evaluate RBTG, and these often spanned different domains. Arora et al. (2024) applied their approach to two projects at Austrian Post, where domain experts confirmed the coherence, feasibility, and comprehensiveness of the generated test scenarios. However, the necessity for human-expert involvement to confirm correctness was also noted.
This challenge highlights the barriers to wider adoption and scaling-up of RBTG methods in industrial settings. Addressing these will require that RBTG approaches align with practical needs, are cost-effective, and can adapt to domain-specific constraints.
11.2. Future work
The challenges identified above should be addressed through more effective and efficient approaches, which should be considered as future RBTG work. This section lists some potential future work for RBTG.
11.2.1. Future Work 1: Exploring LLM-Driven RBTG
A promising future direction will be the application of LLMs to RBTG. Current methods often struggle with ambiguous or incomplete requirements, leading to suboptimal test-case generation. Using the advanced NLP capabilities of LLMs may enable more accurate parsing and clarification of these ambiguous requirements. Once clarified, formal methods could then be applied to generate precise constraints. This would have the potential to significantly enhance the quality and effectiveness of the generated test cases. Further research in this area should also explore how LLMs could be integrated into existing RBTG frameworks, evaluating their performance in real-world industrial settings.
11.2.2. Future Work 2: Optimizing Module-Centric Test Prioritization
Optimizing module-centric test prioritization is another promising future direction. Identifying characteristics of error-prone modules during requirements analysis and test-case design, especially within specific environments, can provide valuable insights and improve testing effectiveness. Since 80% of defects reportedly come from 20% of modules (Boehm and Basili, 2007), the critical modules should be tested more thoroughly.
Although previous research has explored the optimization of test cases through removal of redundant ones (Intana et al., 2023), few studies have adjusted the number of generated test cases based on module criticality. However, it may also sometimes be necessary to tolerate a certain level of redundancy when testing critical modules, to ensure comprehensive testing.
11.2.3. Future Work 3: Engineering Effective RBTG Tools
Despite the abundance of research on RBTG tools, their adoption in industry remains limited. This may be due to a lack of clear evidence that these tools provide significant cost savings or benefits over traditional methods. Developing effective and user-friendly test-generation tools, therefore, is another crucial future direction. This will help bridge the gap between academic research and practical applications, providing industry with useful tools.
11.2.4. Future Work 4: Exploring Non-Functional RBTG
In the reviewed studies, RBTG was predominantly used for functional testing, with limited applications in performance (Ali et al., 2021), security (Chen et al., 2022), and other non-functional requirements. The primary reason for this is that functional requirements constitute the majority of documented requirements. Additionally, performance testing and security testing often require more specialized models that detect and analyze specific types of data (Wang et al., 2024).
Despite the availability of performance tools like LoadRunner (Micro Focus, 2024) and JMeter (Apache Software Foundation, 2024), there remains a need for manual script design. Future work should integrate these tools to automate non-functional testing, reducing manual effort and improving efficiency.
12. Conclusion
This paper has provided a comprehensive survey of 267 requirements-based test generation (RBTG) studies published up to 2024. We traced the distribution of RBTG topics, and analyzed the entire process of generating test cases based on requirements. We classified input requirements’ types and approaches; identified different types of test cases; and conducted a statistical review of existing tools and evaluation methods. We have also highlighted the diverse applications of RBTG across multiple domains, and discussed the challenges and potential opportunities in the field.
Our review revealed that early RBTG methods primarily used formal or semi-formal requirements’ representations to generate test cases, with model-based approaches emerging as the primary strategy. However, recent advancements in natural language processing (NLP) have shifted the focus towards NL requirements. A significant amount of previous research focused on model-based testing, which faces challenges for complex systems. There was also a notable lack of methods for transforming abstract test cases (ATCs) into concrete test cases (CTCs), and a lack of effective evaluation methods.
There are several RBTG challenges that require future investigation. First, input requirements are often not fully integrated with the domain knowledge, the prototype, and other information. Secondly, most test cases generated by RBTG are ATCs, with few universal methods for converting ATCs into CTCs. Finally, current evaluation processes remains insufficient. There is also a lack of general-purpose automated testing tools that can substantially reduce manual effort. We have also identified some potential opportunities, such as the application of large language models (LLMs) to RBTG, potentially enhancing test generation for critical modules.
This paper provides a comprehensive overview of the current state, gaps, and future research directions for RBTG. With automated code generation already widely adopted in industry, the demand for automated testing is more urgent than ever. The insights provided in this paper will benefit both researchers and practitioners, improving testing methodologies and ultimately enhancing the quality and reliability of software systems.
Acknowledgment
This work in this paper was partly supported by the Science and Technology Development Fund of Macau, Macau SAR, under Grant No. 0021/2023/RIA1.
Appendix A List of Surveyed Studies
Table LABEL:tab:listofsurveyedstudies presents the finalized list of 267 research studies surveyed in this work.
Table A.I. List of Surveyed Studies.
No. Title Authors Venue Year
1 Test case generation for specification-based software testing Peng Proceedings of the 1994 Conference of the Centre for Advanced Studies on Collaborative Research 1994
2 A framework for specification-based testing Stocks and Carrington IEEE Transactions on Software Engineering 1996
3 Testing from a Z Specification Hierons RM Software Testing, Verification and Reliability 1997
4 Specification‐based testing using cause‐effect graphs Paradkar et al. Annals of Software Engineering 1997
5 Testing object-oriented programs: from formal specification to test scenario generation Chang et al. Journal of Systems and Software 1998
6 Test scenario generation from a structured requirements specification Cunning and Rozenbiit Proceedings ECBS’99. IEEE Conference and Workshop on Engineering of Computer-Based Systems 1999
7 Using model checking to generate tests from requirements specifications Gargantini and Heitmeyer ACM SIGSOFT Software Engineering Notes 1999
8 A scenario-based approach to validating and testing software systems using statecharts Ryser and Glinz Proceedings of the 12th International Conference on Software and Systems Engineering and their Applications 1999
9 Automated test case generation from dynamic models Fröhlich and Link Proceedings of the European Conference on Object-Oriented Programming 2000
10 UML-Based integration testing Hartmann et al. Proceedings of the 2000 ACM SIGSOFT international symposium on Software testing and analysis 2000
11 Automatic test generation from statecharts using model checking Hong et al. Workshop on Formal Approaches to Testing of Software 2001
12 Synthesis of high-level requirements models for automatic test generation Gupta et al. Proceedings of the 8th Annual IEEE International Conference and Workshop on the Engineering of Computer-Based Systems 2001
13 Test-sequence generation from formal requirement models Rayadurgam and Heimdahl Proceedings of the 6th IEEE International Symposium on High Assurance Systems Engineering 2001
14 Automated generation of statistical test cases from UML state diagrams Chevalley and Thevenod-Fosse Proceedings of the 25th Annual International Computer Software and Applications Conference 2001
15 Requirement-based automated black-box test generation Tahat et al. Proceedings of the 25th annual international computer software and applications conference 2001
16 Generation of functional test sequences from B formal specifications presentation and industrial case-study Legeard and Peureux Proceedings 16th Annual International Conference on Automated Software Engineering 2001
17 Dependence analysis in reduction of requirement based test suites Vaysburg et al. Proceedings of the 2002 ACM SIGSOFT International Symposium on Software Testing and Analysis 2002
18 A UML-based approach to system testing Briand and Labiche Software and Systems Modeling 2002
19 Derivation of domain test scenarios from activity diagrams Reuys et al. Proceedings of the International Workshop on Product Line Engineering 2003
20 A requirement-based approach to test product families Nebut et al. Proceedings of the 2003 International Workshop on Software Product-Family Engineering 2003
21 UML-based statistical test case generation Riebisch et al. Proceedings of the 2002 International Conference on Objects, Components, Architectures, Services, and Applications for a Networked World 2003
22 Generating tests from UML specifications Offutt and Abdurazik Proceedings of the 1999 International Conference on the Unified Modeling Language 2003
23 Scenario-based test case generation for state-based embedded systems Tsai et al. Proceedings of the 2003 IEEE International Performance, Computing, and Communications Conference 2003
24 Automated-generating test case using UML statechart diagrams Kansomkeat and Rivepiboon Proceedings of the 2003 annual research conference of the South African institute of computer scientists and information technologists on Enablement through technology 2003
25 Requirements by contracts allow automated system testing Nebut et al. Proceedings of the 14th International Symposium on Software Reliability Engineering 2003
26 Generating, selecting and prioritizing test cases from specifications with tool support Yu et al. Proceedings of the 3rd International Conference on Quality Software 2003
27 Formal test-case generation for UML statecharts Gnesi et al. Proceedings of the 9th IEEE International Conference on Engineering of Complex Computer Systems 2004
28 Automated functional test case synthesis from THALES industrial requirements Lugato et al. Proceedings of the 10th IEEE Real-Time and Embedded Technology and Applications Symposium 2004
29 RETNA: from requirements to testing in a natural way Boddu et al. Proceedings of the 12th IEEE International Requirements Engineering Conference 2004
30 Generating test cases from UML activity diagram based on Gray-box method Linzhang et al. Proceedings of the 11th Asia-Pacific Software Engineering Conference 2004
31 Model-based system testing of software product families Reuys et al. Advanced Information Systems Engineering 2005
32 UCM-driven testing of web applications Amyot et al. Proceedings of the 12th International SDL Forum on SDL 2005: Model-Driven Engineering for Smart Cities 2005
33 Automating test case generation from Z specifications with Isabelle Helke et al. Proceedings of 10th the International Conference of Z Users 2005
34 Automating the generation and sequencing of test cases from model-based specifications Dick and Faivre International Symposium of Formal Methods Europe 2005
35 Requirements traceability in automated test generation: application to smart card software validation Bouquet et al. ACM SIGSOFT Software Engineering Notes 2005
36 Using adaptive agents to automatically generate test scenarios from the UML activity diagrams Xu et al. Proceedings of the 12th Asia-Pacific Software Engineering Conference 2005
37 Synthesis of scenario based test cases from  B models Satpathy et al. Proceedings of the 2006 International Workshop on Formal Approaches to Software Testing 2006
38 Generating test sequences from UML sequence diagrams and state diagrams Sokenou INFORMATIK 2006–Informatik für Menschen–Band 2, Beiträge der 36. Jahrestagung der Gesellschaft für Informatik eV 2006
39 Automatic test generation: a use case driven approach Nebut et al. IEEE Transactions on Software Engineering 2006
40 Coverage metrics for requirements-based testing Whalen et al. Proceedings of the 2006 international symposium on Software testing and analysis 2006
41 A practical approach for automated test case generation using statecharts Santiago et al. Proceedings of the 30th Annual International Computer Software and Applications Conference 2006
42 Generating interaction test cases for mobile phone systems from use case specifications de Figueiredo et al. ACM SIGSOFT Software Engineering Notes 2006
43 Automatic test case generation from UML communication diagrams Samuel et al. Information and Software Technology 2007
44 Test cases generation from UML activity diagrams Kim et al. Proceedings of the 8th ACIS international conference on software engineering, artificial intelligence, networking, and parallel/distributed computing 2007
45 Test case generation by means of UML sequence diagrams and labeled transition systems Cartaxo et al. Proceedings of the 2007 IEEE International Conference on Systems, Man and Cybernetics 2007
46 An automatic tool for generating test cases from the system’s requirements Ibrahim et al. Proceedings of the 7th IEEE International Conference on Computer and Information Technology 2007
47 Automatic test case generation from UML sequence diagram Sarma et al. Proceedings of the 15th International Conference on Advanced Computing and Communications 2007
48 Automatic test case generation from UML models Sarma and Mall Proceedings of the 10th International Conference on Information Technology 2007
49 Requirement-based testing through formal methods Cabral and Tamai Proceedings of TESTCOM-FATES 2008
50 A novel test case design technique using dynamic slicing of UML sequence diagrams Samuel and Mall e-Informatica Software Engineering Journal 2008
51 Requirements-based test case specification by using information from model construction Pfaller Proceedings of the 3rd International Workshop on Automation of Software Test 2008
52 An approach for supporting system-level test scenarios generation from textual use cases Some and Cheng Proceedings of the 2008 ACM symposium on Applied computing 2008
53 Automatic test case generation using unified modeling language (UML) state diagrams Samuel et al. The Institution of Engineering and Technology 2008
54 Model based testing of system requirements using UML use case models Hasling et al. Proceedings of the 1st International Conference on Software Testing, Verification, and Validation 2008
55 A case study for generating test cases from use cases Gutierrez et al. Proceedings of the 2nd International Conference on Research Challenges in Information Science 2008
56 A transformation-based approach to generating scenario-oriented test cases from UML activity diagrams for concurrent applications Sun Proceedings of the 32nd Annual IEEE International Computer Software and Applications Conference 2008
57 Test Sequence Generation from UML Sequence Diagrams Samuel and Joseph Proceedings of the 9th ACIS International Conference on Software Engineering, Artificial Intelligence, Networking, and Parallel/Distributed Computing 2008
58 Web service test case generation based on decision table (short paper) Noikajana and Suwannasart Proceedings of the 8th International Conference on Quality Software 2008
59 A novel approach to generate test cases from UML activity diagrams Kundu and Samanta Journal of Object Technology 2009
60 Automatic generation of test specifications for coverage of system state transitions Sarma and Mall Information and Software Technology 2009
61 Model-based test cases synthesis using UML interaction diagrams Nayak and Samanta ACM SIGSOFT Software Engineering Notes 2009
62 Test case generation from UML subactivity and activity diagram Fan et al. Proceedings of the 2nd International Symposium on Electronic Commerce and Security 2009
63 Generating test cases for timed systems from controlled natural language specifications Schnelte Proceedings of the 3rd IEEE International Conference on Secure Software Integration and Reliability Improvement 2009
64 TSGen: A UML activity diagram-based test scenario generation tool Sun et al. Proceedings of the 2009 International Conference on Computational Science and Engineering 2009
65 Scenario-based test case generation using event-B models Malik et al. Proceedings of the 1st International Conference on Advances in System Testing and Validation Lifecycle 2009
66 Tracing requirements in a model-based testing approach Abbors et al. Proceedings of the 1st International Conference on Advances in System Testing and Validation Lifecycle 2009
67 Developing comprehensive acceptance tests from use cases and robustness diagrams El-Attar and Miller Requirements Engineering 2009
68 Synthesis of test scenarios using UML activity diagrams Nayak and Samanta Software & Systems Modeling 2009
69 Slicing-based test case generation from UML activity diagrams Samuel and Mall ACM SIGSOFT Software Engineering Notes 2009
70 A hybrid genetic algorithm based test case generation using sequence diagrams Shirole and Kumar Proceedings of the 3rd International Conference on Contemporary Computing 2010
71 Test case generation based on state and activity models Swain et al. Journal of Object Technology 2010
72 Test case generation based on use case and sequence diagram Swain et al. International Journal of Software Engineering 2010
73 Formal scenario-based requirements specification and test case generation in healthcare applications Löffler et al. Proceedings of the 2010 ICSE Workshop on Software Engineering in Health Care 2010
74 Efficient test case generation for validation of UML activity diagrams Chen et al. Design Automation for Embedded Systems 2010
75 Test sequence generation from communicating UML state charts: An industrial application of symbolic transition systems Schwarzl and Peischl Proceedings of the 10th International Conference on Quality Software 2010
76 Automated test case generation from use case: A model based approach Chen and Li Proceedings of the 3rd International Conference on Computer Science and Information Technology 2010
77 Test case generation from Behavioral  UML models Swain and Mohapatra International Journal of Computer Applications 2010
78 Generating test cases from UML activity diagrams using the Condition-Classification Tree Method Kansomkeat et al. Proceedings of the 2nd International conference on software technology and engineering 2010
79 A prolific approach for automated generation of test cases from informal requirements Chatterjee and Johari ACM SIGSOFT Software Engineering Notes 2010
80 A new approach to generating high quality test cases Liu and Miao Proceedings of the 19th IEEE Asian Test Symposium 2010
81 Requirement-based test case generation and prioritization Salem and Hassan Proceedings of the 2010 International Computer Engineering Conference 2010
82 An approach to automatic generation of test cases based on use cases in the requirements phase Kumaran et al. International Journal on Computer Science and Engineering 2011
83 Test Case Generation for Concurrent System using UML Combinational Diagram Khandai et al. International Journal of Computer Science and Information Technologies 2011
84 A proposed test case generation technique based on activity diagrams Boghdady et al. International Journal of Engineering 2011
85 Model based approach to assist test case creation, execution, and maintenance for test automation Gupta and Surve Proceedings of the 1st International Workshop on End-to-End Test Script Engineering 2011
86 Generation of improved test cases from UML state diagram using genetic algorithm Shirole et al. Proceedings of the 4th India Software Engineering Conference 2011
87 A model-based testing using scenarios and constraints-based modular petri nets Reza and Kerlin Proceedings of the 8th International Conference on Information Technology: New Generations 2011
88 A novel approach of test case generation for concurrent systems using UML Sequence Diagram Khandai et al. Proceedings of the 3rd International Conference on Electronics Computer Technology 2011
89 Generating model-based test cases from natural language requirements for space application software Santiago Júnior and Vijaykumar Software Quality Journal 2011
90 Automation of test case generation from textual use cases Jiang and Ding Proceedings of the 4th International Conference on Interaction Sciences 2011
91 An enhanced test case generation technique based on activity diagrams Boghdady et al. Proceedings of the 2011 International Conference on Computer Engineering & Systems 2011
92 Test case generation approach for industrial automation systems Hametner et al. Proceedings of the 5th International Conference on Automation, Robotics and Applications 2011
93 A” Vibration” method for automatically generating test cases based on formal specifications Liu and Nakajima Proceedings of the 18th Asia-Pacific Software Engineering Conference 2011
94 Automatic test case generation from functional requirements in NDT Gutiérrez et al. Proceedings of the 12th International Conference on Web Engineering 2012
95 Test case generation for concurrent object-oriented systems using combinational UML models Dalai et al. International Journal of Advanced Computer Science and Applications 2012
96 Litmus: Generation of test cases from functional requirements in natural language Dwarakanath and Sengupta Proceedings of the 17th International Conference on Applications of Natural Language Processing and Information Systems 2012
97 An enhanced technique for generating hybrid coverage test cases using activity diagrams Boghdady et al. Proceedings of the 8th International Conference on Informatics and Systems 2012
98 Functional and user interface model for generating test cases Bin and Anbao Proceedings of the 11th International Conference on Computer and Information Science 2012
99 Generation test case from UML activity diagram based on AC grammar Pechtanun and Kansomkeat Proceedings of the 2012 International Conference on Computer & Information Science 2012
100 Practical Application of UML activity diagrams for the generation of test cases Fernandez-Sanz and Misra Proceedings of the Romanian Academy, Series A 2012
101 Test generation from state based use case models Nogueira et al. Formal Aspects of Computing 2012
102 Automatic generation of multi-testing types test cases using requirements-based testing Boghdady et al. Proceedings of the 7th International Conference on Computer Engineering & Systems 2012
103 Test case generation based on activity diagram for mobile application Chouhan et al. International Journal of Computer Applications 2012
104 Automatic test cases generation from UML activity diagrams using graph transformation Hettab et al. Proceedings of the 6th International Conference on Information Technology 2013
105 Test case generation from natural language requirements based on SCR specifications Carvalho et al. Proceedings of the 28th Annual ACM Symposium on Applied Computing 2013
106 Extenics-based test case generation for UML activity diagram Li et al. Procedia Computer Science 2013
107 A model based approach to test case generation for testing the navigation behavior of dynamic web applications Bansal and Sabharwal Proceedings of the 6th International Conference on Contemporary Computing 2013
108 Automatic generation of test cases for critical systems based on MC/DC criteria Almeida et al. Proceedings of the IEEE/AIAA 32nd Digital Avionics Systems Conference 2013
109 An approach to generate safety validation test cases from UML activity diagram Tiwari and Gupta Proceedings of the 20th Asia-Pacific Software Engineering Conference 2013
110 Generation of test cases from software requirements using natural language processing Verma and Beg Proceedings of the 6th International Conference on Emerging Trends in Engineering and Technology 2013
111 An automated approach of test case generation for concurrent systems from requirements descriptions Sarmiento et al. Proceedings of the 16th International Conference on Enterprise Information Systems 2014
112 NAT2TESTSCR: Test case generation from natural language requirements based on SCR specifications Carvalho et al. Science of Computer Programming 2014
113 A systematic approach to automatically derive test cases from use cases specified in restricted natural languages Zhang et al. Proceedings of the 2014 International Conference on System Analysis and Modeling 2014
114 Model-based testing from controlled natural language requirements Carvalho et al. Formal Techniques for Safety-Critical Systems 2014
115 A novel approach for test case generation from UML activity diagram Jena et al. Proceedings of the 2014 International Conference on Issues and Challenges in Intelligent Computing Techniques 2014
116 An approach to generate test cases from sequence diagram Dhineshkumar and Galeebathullah Proceedings of the 2014 International Conference on Intelligent Computing Applications 2014
117 Automated generation of test cases from logical specification of software requirements Sharma and Biswas Proceedings of the 9th International Conference on Evaluation of Novel Approaches to Software Engineering 2014
118 Test case generation using UML state diagram and OCL expression AzaharuddinAli et al. International Journal of Computer Applications 2014
119 Generation of test cases using activity diagram Swain et al. International Journal of Computer Science and Informatics 2014
120 Keeping requirements and test cases consistent: Towards an ontology-based approach Feldmann et al. Proceedings of the 12th IEEE International Conference on Industrial Informatics 2014
121 Generation of test cases from sliced sequence diagram Kaur and Singh International Journal of Computer Applications 2014
122 Towards the automated generation of abstract test cases from requirements models Granda et al. Proceedings of the IEEE 1st International Workshop on Requirements Engineering and Testing 2014
123 C&L: Generating model based test cases from natural language requirements descriptions Sarmiento et al. Proceedings of the IEEE 1st International Workshop on Requirements Engineering and Testing 2014
124 Abstract test case generation for behavioural testing of software product lines Devroey et al. Proceedings of the 18th International Software Product Line Conference: Companion Volume for Workshops, Demonstrations and Tools-Volume 2 2014
125 Integration of requirements engineering and test-case generation via OSLC Aichernig et al. Proceedings of the 14th International Conference on Quality Software 2014
126 Cost-effective functional testing of reactive software Venkatesh et al. Proceedings of the 10th International Conference on Evaluation of Novel Approaches to Software Engineering 2015
127 On generating test cases from EDT specifications Venkatesh et al. International Conference on Evaluation of Novel Approaches to Software Engineering 2015
128 Functional test cases generation based on automated generated use case diagram Singh and Sharma International Journal of Innovative Research in Advanced Engineering 2015
129 A model-driven approach for functional test case generation Gutiérrez et al. Journal of Systems and Software 2015
130 Test case generation and optimization using UML models and genetic algorithm Khurana and Chillar Procedia Computer Science 2015
131 A graph transformation approach for automatic test cases generation from UML activity diagrams Hettab et al. Proceedings of the 8th International C* Conference on Computer Science & Software Engineering 2015
132 Automatic generation of system test cases from use case specifications Wang et al. Proceedings of the 2015 International Symposium on Software Testing and Analysis 2015
133 An approach of generating test requirements for agile software development Tiwari and Gupta Proceedings of the 8th India Software Engineering Conference 2015
134 Natural language generation approach for automated generation of test cases from logical specification of requirements Sharma and Biswas Proceedings of the 2014 International Conference on Evaluation of Novel Approaches to Software Engineering 2015
135 Scalable incremental test-case generation from large behavior models Aichernig et al. Proceedings of the 2015 International Conference on Tests and Proofs 2015
136 NAT2TEST tool: From natural language requirements to test cases based on CSP Carvalho et al. Software Engineering and Formal Methods 2015
137 An automated test generation technique for software quality assurance Xu et al. IEEE Transactions on Reliability 2015
138 Model-based approach for automated test case generation from visual requirement specifications Singi et al. Proceedings of the IEEE 8th International Conference on Software Testing, Verification and Validation Workshops 2015
139 Automation of test case generation from behavior tree requirements models Lindsay et al. Proceedings of the 24th Australasian Software Engineering Conference 2015
140 Towards the generation of tests in the test description language from use case map models Boulet et al. Proceedings of the 17th International SDL Forum on SDL 2015: Model-Driven Engineering for Smart Cities 2015
141 Waterfall and agile requirements-based model for automated test cases generation Elghondakly et al. Proceedings of the IEEE 7th International Conference on Intelligent Computing and Information Systems 2015
142 Automatic generation of test cases from activity diagrams for UML based testing (UBT) Oluwagbemi and Asmuni Jurnal Teknologi (Sciences & Engineering) 2015
143 ACT (Abstract to Concrete Tests) - A tool for generating Concrete test cases from formal specification of web applications Bubna and Chakrabarti Proceedings of the 2nd Modelling Symposium 2016
144 A novel technique for generation and optimization of test cases using use case, sequence, activity diagram and genetic algorithm Khurana et al. Journal of Software 2016
145 An optimized approach for automated test case generation and validation for UML diagrams Elghondakly et al. Asian J Inf Technol 2016
146 A transformation-based approach to testing concurrent programs using UML activity diagrams Sun et al. Software: Practice and Experience 2016
147 Automatic test generation from UML sequence diagrams for Android mobiles Anbunathan and Basu International Journal of Applied Engineering Research 2016
148 Automatic generation of specification-based test cases by applying genetic algorithms in reinforcement learning Sato and Sugihara International Workshop on Structured Object-Oriented Formal Language and Method 2016
149 An optimal solution for test case generation using ROBDD graph and PSO algorithm Kalaee and Rafe Quality and Reliability Engineering International 2016
150 Test scenario generation from natural language requirements descriptions based on Petri-Nets Sarmiento et al. Electronic Notes in Theoretical Computer Science 2016
151 Generating test cases from UML activity diagram based on business flow constraints Thanakorncharuwit et al. Proceedings of the 5th International Conference on Network, Communication and Computing 2016
152 Syntactic rules of extracting test cases from software requirements Masuda et al. Proceedings of the 8th International Conference on Information Management and Engineering 2016
153 Test case generation from natural language requirements using CPN simulation Silva et al. Proceedings of the 2015 Brazilian Symposium on Formal Methods 2016
154 Automatic generation of test cases and test purposes from natural language Nogueira et al. Formal Methods: Foundations and Applications 2016
155 Automatic test case generation from requirements for industrial cyber-physical systems Sinha et al. at-Automatisierungstechnik 2016
156 Utilizing business process models to generate software test cases Khader and Yousef International Journal of Computer Science Issues 2016
157 TestMEReq: generating abstract tests for requirements validation Moketar et al. Proceedings of the 3rd International Workshop on Software Engineering Research and Industrial Practice 2016
158 Automatic generation of TestNG tests cases from UML sequence diagrams in Scrum process Elallaoui et al. Proceedings of the 4th IEEE International Colloquium on Information Science and Technology 2016
159 A search-based approach for test suite generation from extended finite state machines Rao et al. Proceedings of the 2016 International Conference on Progress in Informatics and Computing 2016
160 Application of inference rules to a software requirements ontology to generate software test cases Tarasov et al. Proceedings of the 13th International Workshop on Experiences and Directions and 5th International Workshop on Reasoner Evaluation 2017
161 Automated user story driven approach for web based functional testing Masud et al. International Journal of Computer and Information Sciences 2017
162 From statecharts into model checking: A hierarchy-based translation and specification patterns properties to generate test cases de Santiago Júnior  and da Silva Proceedings of the 2nd Brazilian Symposium on Systematic and Automated Software Testing 2017
163 Generating test cases and procedures from use cases in dynamic software product lines Araújo et al. Proceedings of the Symposium on Applied Computing 2017
164 Automated test case generation from UML activity diagram and sequence diagram using depth first search algorithm Meiliana et al. Procedia Computer Science 2017
165 Constructing test cases using natural language processing Ansari et al. Proceedings of the 3rd International Conference on Advances in Electrical, Electronics, Information, Communication and Bio-Informatics 2017
166 Transforming workflow models into automated end-to-end acceptance test cases Boucher and Mussbacher Proceedings of the IEEE/ACM 9th International Workshop on Modelling in Software Engineering 2017
167 Automated test case generation from high-level logic requirements using model transformation techniques Olajubu et al. Proceedings of the 9th Computer Science and Electronic Engineering 2017
168 Towards the generation of end-to-end web test scripts from requirements specifications Clerissi et al. Proceedings of the IEEE 25th International Requirements Engineering Conference Workshops 2017
169 Reqcap: Hierarchical requirements modeling and test generation for industrial control systems Almohammad et al. Proceedings of the IEEE 25th International Requirements Engineering Conference Workshops 2017
170 An automated test case generation approach based on activity diagrams of SysML Yin et al. International Journal of Performability Engineering 2017
171 Automatic test case generation through collaboration diagram: a case study Kaur and Vig International Journal of System Assurance Engineering and Management 2017
172 Requirement-based testing-extracting logical test cases from requirement documents Sneed Software Quality: Methods and Tools for Better Software and Systems 2017
173 An automated functional testing framework for context-aware applications Mirza and Khan IEEE Access 2018
174 Named entity recognition for automated test case generation Mahalakshmi et al. The International Arab Journal of Information Technolog 2018
175 Automatic test case generation method for large scale communication node software Kikuma et al. Proceedings of the 2018 International Conference on Emerging Internetworking, Data & Web Technologies 2018
176 Automated generation of requirements-based test cases for an adaptive cruise control system Aniculaesei et al. 2018 IEEE Workshop on Validation, Analysis and Evolution of Software Tests 2018
177 Automatic generation of test cases from  UML models Pérez  and Marín CLEI Electronic Journal 2018
178 Automatic test transition paths generation approach from EFSM using state tree Chen et al. Proceedings of the 2018 IEEE International Conference on Software Quality, Reliability and Security Companion 2018
179 From natural language requirements to passive test cases using guarded assertions Danie et al. Proceedings of the 2018 IEEE International Conference on Software Quality, Reliability and Security 2018
180 Automatic test cases generation from business process models Yazdani et al. Requirements Engineering 2018
181 A transformation-based method for test case automatic generation from use cases Hue et al. Proceedings of the 10th International Conference on Knowledge and Systems Engineering 2018
182 Test scenarios generation using UML sequence diagram Panigrahi et al. Proceedings of the 2018 International Conference on Information Technology 2018
183 From modeling to test case generation in the industrial embedded system domain Hussain et al. Proceedings of the 2018 International Conferences on Software Technologies: Applications and Foundations 2018
184 Automated generation of requirements-based test cases for an automotive function using the SCADE toolchain Aniculaesei et al. Proceedings of the 11th International Conference on Adaptive and Self-Adaptive Systems and Applications 2019
185 An efficient method to generate test cases from UML-use case diagram Khalifa and Jamil International Journal of Engineering Research and Technology 2019
186 From requirements to automated acceptance tests of interactive apps: An integrated model-based testing approach Maciel et al. Proceedings of the 14th International Conference on Evaluation of Novel Approaches to Software Engineering 2019
187 CPN simulation-based test case generation from controlled natural-language requirements Silva et al. Science of Computer Programming 2019
188 A memorization approach for test case generation in concurrent UML activity diagram Kamonsantiroj et al. Proceedings of the 2019 2nd International Conference on Geoinformatics and Data Analysis 2019
189 Ontology based test case generation for black box testing Ul Haq and Qamar Proceedings of the 2019 8th International Conference on Educational and Information Technology 2019
190 From use case maps to executable test procedures: a scenario-based approach Kesserwan et al. Software & Systems Modeling 2019
191 Using OWL ontologies as a domain-specific language for capturing requirements for formal analysis and test case generation Crapo and Moitra Proceedings of the IEEE 13th International Conference on Semantic Computing 2019
192 MCCFG: an MOF-based multiple condition control flow graph for automatic test case generation Son et al. Cluster Computing 2019
193 Test case generation using unified modeling language Shah et al. Proceedings of the International Conference on Computer and Information Sciences 2019
194 An automated approach to generate test cases from use case description model Alrawashed et al. Computer Modeling in Engineering & Sciences 2019
195 Specification-based test case generation with genetic algorithm Wang et al. Proceedings of the 2019 IEEE Congress on Evolutionary Computation 2019
196 Automating requirements analysis and test case generation Moitra et al. Requirements Engineering 2019
197 Towards transforming user requirements to test cases using MDE and NLP Allala et al. Proceedings of the IEEE 43rd Annual Computer Software and Applications Conference 2019
198 Test case generation, selection and coverage from natural language Nogueira et al. Science of Computer Programming 2019
199 Using the SCADE toolchain to generate requirements-based test cases for an adaptive cruise control system Aniculaesei et al. Proceedings of the ACM/IEEE 22nd International Conference on Model Driven Engineering Languages and Systems Companion 2019
200 USLTG: Test Case Automatic Generation by Transforming Use Cases Hue et al. International Journal of Software Engineering and Knowledge Engineering 2019
201 Requirements-based automated test generation for safety critical software Li et al. Proceedings of the 38th Digital Avionics Systems Conference 2019
202 Generating test sequences from UML use-case diagrams Hamza and Hammad Proceedings of the 2019 International Conference on Innovation and Intelligence for Informatics, Computing, and Technologies 2019
203 Enhancing software testing with ontology engineering approach Charoenreh and Intana Proceedings of the 23rd International Computer Science and Engineering Conference 2019
204 Impact analysis framework of test cases based on changes of use case based requirements Intana and Sriraksa Proceedings of the 23rd International Computer Science and Engineering Conference 2019
205 Knowledge graph based automated generation of test cases in software engineering Nayak et al. Proceedings of the 7th ACM IKDD CoDS and 25th COMAD 2020
206 Test case generation algorithms and tools for specifications in natural language Aoyama et al. Proceedings of the 2020 IEEE International Conference on Consumer Electronics 2020
207 From requirements to automated acceptance tests with the RSL language Paiva et al. Proceedings of the 2019 International Conference on Evaluation of Novel Approaches to Software Engineering 2020
208 Automatic generation and optimization of test case using hybrid cuckoo search and bee colony algorithm Lakshminarayana and SureshKumar Journal of Intelligent Systems 2020
209 Model-based testing from model driven architecture: A novel approach for automatic test cases generation Essebaa et al. Proceedings of the 2019 International Conference on Smart Information & Communication Technologies 2020
210 Preliminary experiences in requirements-based security testing Miranda et al. Proceedings of the 2020 International Conference on the Quality of Information and Communications Technology 2020
211 Scaling test case generation for expressive decision tables Agrawa et al. Proceedings of the IEEE 13th International Conference on Software Testing, Validation and Verification 2020
212 SPECMATE: Automated creation of test cases from acceptance criteria Fischbach et al. Proceedings of the IEEE 13th International Conference on Software Testing, Validation and Verification 2020
213 SYNTest: Prototype of syntax test case generation tool Intana et al. Proceedings of the 5th International Conference on Information Technology 2020
214 Automatic test case generation from formal requirement model for avionics software Wang et al. Proceedings of the 6th International Symposium on System and Software Reliability 2020
215 Test generation tool for modified condition/decision coverage: model based Testing El qortobi et al. Proceedings of the 13th International Conference on Intelligent Systems: Theories and Applications 2020
216 Specification-based test case generation with constrained genetic programming Sato Proceedings of the 20th International Conference on Software Quality, Reliability and Security Companion 2020
217 NLP-based requirements formalization for automatic test case generation Gröpler et al. Proceedings of the 29th International Workshop on Concurrency, Specification and Programming 2021
218 Towards a test case generation tool based on functional requirements Fernandes and Maciel Proceedings of the 19th Brazilian Symposium on Software Quality 2021
219 Test case generation from UML-diagrams using genetic algorithm Sahoo et al. Computers, Materials and Continua 2021
220 Towards a model-driven testing framework for GUI test cases generation from user stories Granda et al. Proceedings of the 16th International Conference on Evaluation of Novel Approaches to Software Engineering 2021
221 Analyzing UML use cases to generate test sequences Hamza and Hammad International Journal of Computing and Digital Systems 2021
222 Executable test case generation from specifications written in natural language and test execution environment Aoyama et al. Proceedings of the IEEE 18th Annual Consumer Communications & Networking Conference 2021
223 An ontology-based approach to automated test case generation Banerjee et al. SN Computer Science 2021
224 Accuracy improvement by training data selection in automatic test cases generation method Ueda and Tsukada Proceedings of the 9th International Conference on Information and Education Technology 2021
225 Model-based test case generation from UML sequence diagrams using extended finite state machines Rocha et al. Software Quality Journal 2021
226 A method of test case generation based on VRM model Zhipeng et al. Proceedings of the IEEE 6th International Conference on Computer and Communication Systems 2021
227 Syntax-tree similarity for test-case derivability in software requirements Masuda et al. Proceedings of the 2021 IEEE International Conference on Software Testing, Verification and Validation Workshops 2021
228 Coverage-based grammar-guided genetic programming generation of test suites Ibias et al. Proceedings of the 2021 IEEE Congress on Evolutionary Computation 2021
229 An improved approach for generating test cases during model-based testing using tree traversal algorithm Achimugu et al. Journal of Software Engineering and Applications 2021
230 Automated test cases generation from requirements specification Lafi et al. Proceedings of the 2021 International Conference on Information Technology 2021
231 Generating test scenarios using SysML activity diagram Yang et al. Proceedings of the 8th International Conference on Dependable Systems and Their Applications 2021
232 Generating test cases from requirements: A case study in railway control system domain Zheng et al. Proceedings of the 2021 International Symposium on Theoretical Aspects of Software Engineering 2021
233 Automatic generation of test cases from formal specifications using mutation testing Cajica et al. Proceedings of the 18th International Conference on Electrical Engineering, Computing Science and Automatic Control 2021
234 Model-based test case generation approach for mobile applications load testing using OCL enhanced activity diagrams Ali et al. Proceedings of the 10th International Conference on Intelligent Computing and Information Systems 2021
235 A tool to support vibration testing method for automatic test case generation and test result analysis Saiki et al. Proceedings of the IEEE 21st International Conference on Software Quality, Reliability and Security 2021
236 Forward-propagation approach for generating feasible and minimum test case suites from cause-effect graph specifications Krupalija et al. IEEE Access 2022
237 Application of the equivalent partitioning method in testing for automatic test case generation on the Digi-OTA system Amalia et al. Jurnal Media Informatika Budidarma 2022
238 Automatic generation of acceptance test cases from use case specifications: An NLP-based approach Wang et al. IEEE Transactions on Software Engineering 2022
239 Automatic test case and test oracle generation based on functional scenarios in formal specifications for conformance testing Liu and Nakajima IEEE Transactions on Software Engineering 2022
240 Transition coverage based test case generation from state chart diagram Pradhan et al. Journal of King Saud University-Computer and Information Sciences 2022
241 Automating test case generation for android applications using model-based testing Khan et al. EMITTER International Journal of Engineering Technology 2022
242 Prioritized variable-length test cases generation for finite state machines Rechtberger et al. Proceedings of the 2022 IEEE International Conference on Software Testing, Verification and Validation Workshops 2022
243 Requirements-driven model checking and test generation for comprehensive verification Bhatt et al. Proceedings of the 14th International Symposium NASA Formal Methods 2022
244 Model-based automated testing method for airborne system Zhu et al. Proceedings of the 5th International Conference on Aerospace System Science and Engineering 2022
245 MC/DC test case automatic generation for safety-critical systems Wang et al. Proceedings of the IEEE 22nd International Conference on Software Quality, Reliability and Security 2022
246 An improved test case generation method based on test requirements for testing software component Chen et al. Proceedings of the IEEE 22nd International Conference on Software Quality, Reliability, and Security Companion 2022
247 Generating abstract test cases from user requirements using MDSE and NLP Allala et al. Proceedings of the IEEE 22nd International Conference on Software Quality, Reliability and Security 2022
248 Model-checking-based automated test case generation for Z formal specification of an urban railway interlocking system Kadakolmath and Ramu Proceedings of the 4th International Conference on Emerging Research in Electronics, Computer Science and Technology 2022
249 Automatic creation of acceptance tests by extracting conditionals from requirements: NLP approach and case study Fischbach et al. Journal of Systems and Software 2023
250 Automated test case generation using T5 and GPT-3 Mathur et al. Proceedings of the 9th International Conference on Advanced Computing and Communication Systems 2023
251 Generating robot framework test scripts from user stories and scenarios for web application testing Waitchasarn et al. Proceedings of the 8th International Conference on Computer and Communication Systems 2023
252 Research on software test case generation based on SysML Chen et al. Proceedings of the 2023 International Symposium on Intelligent Robotics and Systems 2023
253 Decision tree based test case generation using story board and natural language processing Gupta et al. Proceedings of the 2023 International Conference on Advances in Computing and Data Sciences 2023
254 Transforming software requirements into test cases via model transformation Kesserwan et al. International Journal of Software Engineering & Applications 2023
255 An approach of test case generation with software requirement ontology Intana et al. International Journal of Advanced Computer Science and Applications 2023
256 Granular traceability between requirements and test cases for safety-critical software systems Elqortobi et al. Proceedings of the 19th International Conference on Mobile Web and Intelligent Information Systems 2023
257 Automatic test sequence generation and functional coverage measurement from UML sequence diagrams Ekici and Tuglular International Journal of Information System Modeling and Design 2023
258 Test case generation and history data analysis during optimization in regression testing: An NLP study Gupta and Mahapatra Cogent Engineering 2023
259 Generating automated test case from sequence diagram using Pre-order Traversal Yonathan et al. Procedia Computer Science 2024
260 Test case information extraction from requirements specifications using NLP-based unified boilerplate approach Lim et al. Journal of Systems and Software 2024
261 A method for selecting training data using Doc2Vec for automatic test Cases generation Fujita and Ueda Proceedings of the 2024 IEEE International Conference on Consumer Electronics 2024
262 A case study: To discuss the construction of the finite state machine and markov-chain OP models for test case generation Shaheen et al. Proceedings of the 5th International Conference on Advancements in Computational Sciences 2024
263 Generating test scenarios from NL requirements using retrieval-augmented LLMs: An industrial study Arora et al. Proceedings of the IEEE 32nd International Requirements Engineering Conference 2024
264 Automatic test cases generation with selection of training data for various system specifications Ueda et al. Proceedings of the 6th International Conference on Computer Communication and the Internet 2024
265 User story based automated test case generation using NLP Chinnaswamy et al. Proceedings of the 2024 International Conference on Computational Intelligence in Data Science 2024
266 Automated test case generation for satellite FRD using NLP and large language model Shakthi et al. Proceedings of the 4th International Conference on Electrical, Computer, Communications and Mechatronics Engineering 2024
267 Leveraging pre-trained large language models (LLMs) for on-premises comprehensive automated test case generation: An empirical study Yin et al. Proceedings of the 9th International Conference on Intelligent Informatics and Biomedical Sciences 2024